### v1

# The Chemistry and Physics of 199Hg Nuclear Spin Polarization Relaxation in Quantum Magnetometry Cells

Abstract
Quantum spin magnetometry using optically pumped 199Hg has been successfully used in many fundamental physics experiments. A serious problem that has not been resolved is the instability of the 199Hg spin relaxation rate in atomic vapor cells under irradiation with 254 nm Hg resonance light. In this paper, previously obtained data are re-analyzed or analyzed for the first time. The effects of impurities of H2 and O2 are elucidated, and possible ways to stabilize cells are discussed. Surface states originating from the an der Waals interaction of 199Hg with fused silica are analyzed and shown to be critical to understanding relaxation mechanisms. A discussion of the possible use of a mixture of N2O and other gases is presented.
quantum spin magnetometry; optical pumping; photosensitized reactions; van der Waals interactions.
I Introduction
This research article is different from most because it describes work using optically pumped 199Hg that has been done over the last 40 years, some of which was never published, and with only partial theoretical analysis. As such, a review article is not possible due to the limited specific literature to review. Typically, studies of cell atomic polarization relaxation rates (or inverse of the spin lifetime) were done in the context of building an apparatus to measure some fundamental process of greater immediate interest; as typical in such work, the tedious technical details are barely left as a footnote if presented at all.
In this research article, optically pumped 199Hg will be put in context with other work in the field. In particular, the importance of surface states and the van der Waals interaction of atoms with walls will be discussed together with a theoretical analysis.
The optical pumping of the ground states of 199Hg and 201Hg was first achieved in the early 1960’s by B. Cagnac at the Laboratoire de Spectroscopie Hertzienne de l’ENS (now named the Kastler-Brossel Laboratory). [1] The study of the optical pumping of ground state atoms had been expanding through the 1950’s and Hg was among the very atoms first to achieve spin relaxation times of 100’s of seconds. (See [2, 3] for a review of the field.) Hg is a 1S0 atom, so the ground-state magnetic moment is due only to the nucleus; for 199 $I=1/2$, and for 201, $I=3/2$. Other than application to nuclear magnetic resonance-based gyroscopes [4, 5], optically pumped Hg has found minimal practical applications because alkali atoms, with an unpaired electron spin, provide much greater sensitivities in magnetometer applications.
In the early 1980’s at the University of Washington, Seattle, Prof. E.N. Fortson identified optically pumped Hg as useful for a number of fundamental experiments, the first to search for spatial anisotropy that used the mass quadupole moment of the 201Hg nuclear spin to limit the coupling to its orientation or motion through absolute space in a modernized Hughes-Drever experiment.[6, 7]. The optical pumping cells developed for this work were also used to test a proposed non-linear extension to quantum mechanics, which incidentally provided a gyroscopic measurement of the Earth’s rotation.[8]
These experiments used optical pumping cells that were heated to around 400 C and used the 199Hg nuclear spin as a comagnetometer (a term we invented, using “co" from Latin, meaning “with" or “together") to eliminate frequency changes due to magnetic field fluctuations that would interfere with the fundamental physics measurements via 201Hg.
However, the main goal of that work was to develop an optical pumping cell in which an electric field could be applied to the processing of 199Hg atoms. If the atom has an electric dipole moment in addition to the magnetic one, there would be a shift in the Larmor frequency and would imply the existence of a fundamental time-reversal violating interaction(s) with the 199 Hg atom, the effects of which are expected to be enhanced in high atomic number atoms. (See [9] for a review of the field.)
The goal at the time (1983-5) was to find a cell configuration that would give a long 199Hg spin polarization lifetime and allow the application of a strong electric field to the atoms, without the need to heat the cell to high temperature and also with a buffer gas (both for high voltage stability). An earlier experiment in the Fortson group employed spin-exchange polarized 129Xe,[10] for an EDM experiment, and lifetimes near room temperature on the order of a thousand or more seconds were routinely achieved in the presence of a field on the order of a few kV/cm. As 199Hg offers a much higher sensitivity to possible fundamental interactions of interest in addition to intrinsically better measurement sensitivity, a vigorous research effort was launched. Attempts to find another spin-1/2 atom to serve as a comagnetometer in the same cell as 199Hg were unsuccessful; however, a workable cell was developed with 199Hg alone, although the spin polarization lifetime would degrade with exposure to 254 nm Hg resonance radiation. This problem persists to this day in that the construction of stable cells appears as hit-and-miss.
The purpose of this paper is to briefly review the development of 199Hg optical pumping cells for EDM and other fundamental experiments and to provide an analysis of data that has not been published. This data sheds light on the chemical and physical mechanisms of lifetime stability, which, when elucidated, will help identify what is needed to attain reproducibility in cell fabrication.
Optically pumped 199Hg has found use in several other fundamental experiments, for example, a tests for axions and their long-range interaction from both spin-polarized and unpolarized material bodies.[11, 12] All of these experiments, including the 199Hg EDM experiments, employ closed cells in that they are filled with appropriate amounts of 199Hg and other gases and permanently sealed off from a preparation manifold.
In addition, 199Hg has served as a comagnetometer for several neutron EDM experiments, the first application being at the Institut Laue-Langevin.[13, 14] In these experiments, the storage cell for ultracold neutrons (UCN) and 199Hg together needed to be developed, and lifetime stability issues persisted in these experiments. A coating that is suitable for both the atomic spin and storage of UCN was found [15], and to date, there does not yet appear to be a fully satisfactory replacement. These experiments are “open" in that background gases (including 199Hg) can be pumped out of the cell.
The spin relaxation lifetime issues for the closed and open cells are related, as will be examined here. There has been continued interest in neutron EDM and specifically experiments that employ a 199Hg comagnetometer.[16] Although the 199Hg EDM limit[17] is numerically much lower than the neutron EDM (this allows it to be used as a comagnetometer), interpretation of the former results in a sensitivity loss regarding fundamental interactions, so the impacts of both types of experiments on testing the standard model of electroweak interactions are similar. New neutron EDM experiments employing 199Hg magnetometry are currently being developed.[18, 19]
I.1 Development of 199Hg EDM cells
Research toward developing a room temperature 199Hg optical pumping cell was begun in 1984, with the first trials using blown fused-silica bulbs of about 18 mm ID. The kinetic mean free path (mfp) between wall collisions is given by
$\ell=\frac{4V}{A}=\frac{4}{3}R$ (1)
where $V$ and $A$ are the internal volume and surface area, and $R$ is the sphere radius. These cells also had fill stems about 20 mm long where they were sealed from the vacuum system after being filled with natural Hg 0 C vapor pressure.
The first cell had the inner surface coated with paraffin (dotriacontane) and had a spin polarization lifetime of about 40 s. For a larger cell as envisioned for EDM work, with a factor of 2 greater mfp, assuming homogeneous surface relaxation, the lifetime would be 80 s, and good enough for an EDM experiment.
Under illumination with the 254 nm pumping light, the cell lifetime deteriorated. It could be temporarily restored by remelting the wax; overall, ominous effects that we did not fully appreciate at the time.
The next bath of spherical cells had siliconizing agents applied to the inner surface, both Surfasil® and Aquasil® were tested, both with and without wax, and one Aquasil® cell was filled with 40 torr N2 buffer gas (Linde® spectroscopic grade, of unknown history). The wax was introduced only after the cells without wax were removed from the vacuum system. It was observed that the wax did not wet the siliconized surfaces but accumulated in individual droplets, so the wax was driven into the cell stem, where it formed a sort of crystalline plug.
All of these cells had a lifetime of only about 15 sec. This was thought to be due to the damage to the coating that occurred with the cells being sealed off, the poor character of the wax plug, or magnetic field gradients in the case of the buffer gas cell (the magnetic shield was very crude), which had a 5 s. lifetime. The Aquasil® coated cell (10% dilution in acetone, followed by two rinses with acetone and drying by blowing air into the bulb) had the best long-term performance. Warming the cell to 120 $\circ$ C increased useful life to 60 s, with a modest permanent improvement at room temperature.
Emboldened by these results, we made cells similar to those used in [10] but with Hg and only N2 as a buffer gas that is required for high-voltage stability. These cells had a 25 mm OD, 23 mm ID, 10 cm tall fused silica cylinder with a fill stem that enters radially into the side of the cylinder, and capped by fused silica disks of approximately 38mm diameter, glued on with Varian Torr-Seal® epoxy. For Hg, a variety of electrode materials were tried; however, exactly the same electrode coating material (tin oxide) coated with wax with Aquasil® on the cylinder body gave excellent performance.
In Fig. 1, a free-precession signal is shown, obtained by optical pumping with resonance light from a 204Hg microwave discharge lame; there is no hyperfine structure for this isotope, and the single spectral line overlaps the $F=1/2$ to $F=1/2$ level for 199Hg, and the $F=3/2$ to $F=5/2$ for 201Hg (for which no optical pumping signal was ever observed in our EDM cells, as expected).
This plot was obtained by optically pumping the system with circularly polarized resonance light propagating along the direction of a weak magnetic field (10 mG), which was suddenly switched off with the simultaneous sudden application of a second field (10 mG) perpendicular to the light. The angular momentum states for the transition in 199Hg are optimal for optical pumping because the sample becomes transparent (for pure 199Hg) when fully polarized, and the modulation on the light during free precession is 100%, however that modulation is reduced to about 15% with a natural isotopic mixture of Hg.(See Fig. 1 of [2] for a pictorial of this specific case.)
Remarkably, the relaxation time in the dark was about 3600 s. We therefore thought that we were done with cell development, as cells made in this manner proved to be stable under the application of high voltage.
Figure 1: 1st successful EDM cell, probe light limited lifetime to 200 sec, dark lifetime was 3600 s at the time of the measurement.
Unfortunately, our hopes were dashed when we discovered that the cells were unstable against the loss of Hg. The first version of the 199Hg experiment was based on an atomic oscillator configuration, where the static magnetic field (quantization axis) is oriented 45 away from the light propagation direction, and an oscillating field close to the Larmor frequency is applied. The atomic system develops a coherent polarization oscillation, with the precession evident in the transmitted light modulation. The phase shift between the atomic precession and the oscillating field is proportional to the degree of detuning between the Larmor and oscillating field frequencies and can be measured with a phase-sensitive detector.
Two such cells were operated together with electric fields directed oppositely, with the sum of the frequency difference used to stabilize the applied field and a correction signal to stabilize the gradient field and which provided a signal to detect an EDM because the field in the cells was directed oppositely.[20]
The in-phase transmitted light signal provides a measure of the sample atomic polarization, times the atomic density. Shown in Fig. 2 is one of the data runs where the loss of Hg one of the cells is evident as the average transmitted light intensity, measured by a solar blind photomultiplier, steadily increases, representing a loss of Hg atoms (the discharge lamps used in this work were remarkably stable, with few percent at most change in 24 hours). At the same time, the in-phase signal substantially increases, and this is only possible if the wall lifetime increases, which increases the atomic polarization for the approximately constant optical pumping rate.
Figure 2: $V_{DC}$ is a measure of the transmitted light intensity (multiplied by 10 to offset the plot from $V_{in}$), and its increase indicates a loss of Hg atoms. The atomic oscillator in phase signal continues to increase, which is only possible if the cell relaxation rate decreases (e.g. lifetime increases), and then disappears when the Hg density falls toward zero. Data manually digitized from plots in [7], Fig. 8.2, cell 2.
For the case of low absorption, the transmitted light is, by expanding $e^{-x}\approx 1-x$,
$I(t)=I_{0}\left[1-\rho(t)\sigma\ell\big{(}1-\epsilon P(t)\cos\omega t\big{)}\right]$ (2)
where $I_{0}$ is the resonant light intensity, $\rho(t)$ is the time-dependent Hg density, $\sigma$ is the light absorption cross section, $\ell$ is the effective path length for light through the cell, $\epsilon$ is the fraction of light resonant with 199Hg , $P(t)$ is the average magnitude of the atomic polarization, and $\omega$ is the Larmor frequency. Therefore the average $DC$ light intensity, represented by $V_{DC}(t)$, is
$V_{DC}(t)\propto I_{0}\big{(}1-\rho(t)\sigma\ell\big{)}$ (3)
and the in-phase signal is
$V_{in}\propto I_{0}\rho(t)\sigma\ell P(t).$ (4)
The fractional change in the amount of light absorbed by the atoms, relative to the initial, is
$\eta(t)=\frac{V_{DC}(\infty)-V_{DC}(t)}{V_{DC}(\infty)-V_{DC}(0)}$ (5)
$V_{dc}(\infty)\approx V_{dc}(t_{f})+30$ (6)
in the voltage units of Fig. 2, and where $t_{f}=18000$ s is the end of the measurement time however the absorption has not yet reached zero. A rough extrapolation gives a small increase in the absorption.
It is the absorbed light that produces $V_{in}$, so
$\frac{V_{in}(t)}{\eta(t)}\propto P(t).$ (7)
From the basic principles of the atomic oscillator,
$P(t)=\frac{\Gamma_{p}}{\Gamma_{p}+\Gamma_{w}+\omega_{1}}$ (8)
where $\Gamma_{p}$ is the optical pumping rate, $\Gamma_{w}=1/\tau_{w}$ is the spin relaxation rate and $\omega_{1}$ is the effective precession rate around the oscillating field.
Because the in-phase atomic oscillator signal $V_{in}(t)$ is proportional to the absorbed light $\eta(t)$, it follows that
$\frac{\eta(t)}{V_{in}(t)}\propto(P(t))^{-1}=\frac{\Gamma_{p}(t)+\Gamma_{w}(t)+\omega_{1}}{\Gamma_{p}(t)}.$ (9)
The pumping rate is given by the average DC light intensity in the cell, which is
$\Gamma_{p}\propto[1-\alpha\eta(t)]$ (10)
where $\alpha=0.58$ is the initial light absorption as determined by $V_{dc}(0)$ and $V_{dc}(\infty)$, noting that typically 20% of the light is nonresonant for the microwave discharge lamps we employed.
Therefore,
$\frac{[1-\alpha\eta(t)]\eta(t)}{V_{in}(t)}\propto\Gamma_{p}(t)+\Gamma_{w}(t)+\omega_{1}=\Gamma_{tot}(t).$ (11)
We can take the ratio $\Gamma_{tot}(t)/\Gamma_{tot}(0)$ to determine how the total relaxation lifetime evolves, which is shown in Fig. 3 to decrease. This can be compared to the pumping rate $\Gamma_{p}(t)/\Gamma_{p}(0)$, which increases over the same time period, because the light intensity in the cell increases as Hg is lost. The other term in $\Gamma_{tot}(t)$ is $\omega_{1}$ which is constant.
In typical operation of the atomic oscillators, the light intensity and $\omega_{1}$ were kept constant between data runs. Normally we tried to achieve
$\omega_{1}=\Gamma_{p}\approx\Gamma_{w}/2.$ (12)
This relationship can only be satisfied for the data in Fig. 3 if $\Gamma_{w}(0)>5\Gamma_{p}$, a factor 2.5 times larger than the target value. However, this choice provided a reasonable balance over the run, plus stronger pumping tended to degrade the cells faster.
This shows that, without a doubt, during this data run, $\Gamma_{tot}(t)$ decreased with time, so $\tau_{w}$ must have increased as Hg was consumed. From this graph, we estimate that $\tau_{w}$ increased by at least a factor of three over the duration of the experimental run (data near the end of the run are unreliable due to the failure of the atomic oscillator as the signal diminished) to within the approximations and the lack of experimental specifics for work done nearly 40 years ago. However, this is the first clear evidence that 254 nm resonance radiation irradiation can be beneficial in one way, and perhaps the only direct evidence to date.
In principle, a source of Hg in the cell could be along with a small amount of oxygen; however, the optical density would be too high for any practical system at the time.
Figure 3: By use of the cell light transmission, it is possible to deduce the the pump rate $\Gamma_{p}$, and the in-phase signal to deduce $\Gamma_{p}+\Gamma_{w}+\omega_{1}$. As the last term is constant, and their sum is decreasing while $\Gamma_{p}$ is increasing, is clear that the wall relaxation rate is decreasing. The loss of Hg atoms eventually lead to a failure of the atomic oscillator for this cell, near the final time of the plotted data.
In obtaining the first 199Hg EDM limit, new cells were manufactured for each short data run and the result of that was published [7, 20]. .
Table 1: Excitation and bond energies of interest. Most are available in chemistry textbooks (e.g., [21]); citations are provided for the rarely encountered bonds.
State Energy Comment/Reference
or Bond (enthalpy) $[e{\rm V}]$
3P1 Hg 4.89 allowed transition ($f\approx 1$)
3P0 Hg 4.55 metastable, populated by N2 collisions with 3P1 Hg
H-H 4.45
O-H 4.76
(H:O)-H 4.8
C-H 4.26
F-H 5.80
Hg-H 0.37 the extra electron in HgH destabilizes the bond
(H:Hg)-H 3.0 thermodynamically unstable [22]
O-Hg 0.94
C-Hg 0.72 [23]
(O:O):O 1.5
C-F 4.52
C-C 3.58
C$=$O 8.23
C$=$C 8.62
O$=$O 5.12
(O:C)-H 1.64 formyl (free) radical,
0.087 $e$V activation energy to form formaldehyde[24]
(HCO)-H 3.82 formaldehyde
[…]
(N2):O -0.85 exothermic, 1.7 $e$V activation energy (can detonate) [25];
scavenges O3 [26]
I.2 Oxygen
The next stage of experimental development was to identify the Hg loss mechanism. A chapter in [27] describes photosensitized reactions involving 6 3P1 excited-state Hg atoms. Specifically, excited Hg in collisions with O2 transfers most of its energy to the molecule, which subsequently reacts with another O2 to form O3 (ozone) and atomic oxygen O. O3 subsequently reacts with Hg to form HgO and O2, thus losing an Hg atom from the gas. There is no isotopic enrichment in the HgO formed in this process, which led to the full elucidation of this process [28] that has been known for close to 100 years.[29] It can be understood that such a series of reactions is possible from the bond energies given in Table 1.
Armed with this knowledge, we took steps to remove oxygen from the nitrogen buffer gas by bubbling the gas through liquid amalgams of Mg-Hg or Al-Hg followed by a liquid nitrogen trap to remove Hg vapor.[30] After such purification, the Hg density became constant; however, the cell lifetimes would rapidly degrade with resonance light irradiation.
A further study of photosensitized reactions indicated that metastable 6 3P0 atoms, which are readily formed with collisions between 6 3P1 atoms and nitrogen molecules, due to resonance with the rotation states of N2. The 6 3P0 excited atoms still carry considerable energy, see Table 1, and can damage the wall coatings. The metastable lifetime has been estimated as seconds, so there is ample time for diffusion to the cell wall.
The destructive nature of 6 3P1 and 6 3P0 atoms has been known for nearly 100 years (this was known before the possibility of producing ozone [31]). As shown in Table 1, these excited states have enough energy to break most single bonds.
Indeed, even bare fused silica is prone to damage by these excited atoms. In our earlier work with testing spatial isotropy, we discovered that fused silica optical pumping cells that were irradiated while at room temperature would suffer irreversible damage as manifested by a shorted spin lifetime. The damaging effects have been known since 1930 [32] but apparently not by us.
There was a vague notion that hydrogen was responsible for the degradation of the spin lifetime; however, a model was never developed, and the notion tended to be dismissed because of the successful operation of hydrogen masers (more on this later).
The best option to quench 6 3P1 (that is, perturb the excited state causing a radiative transition or energy transfer) directly to the ground state and also to quench 6 3P0 (which is not quenched by nitrogen) is carbon monoxide, CO. Cells were made with 5% CO and 95%N2 and the cell life expectancy improved. UW 199Hg EDM experiments over the next decade employed such cells.
I.3 Further Advances
The latest 199Hg EDM experiment is reported in [17], and incorporated a number of improvements. First, it used a pulsed pump and probe method that gave a factor of nearly 2 in the average atomic polarization, as suggested in [33]. This was demonstrated together with spin precession detected by the Faraday rotation of the off-resonant light provided by an 200Hg discharge lamp. (It should be noted that a similar system is described in [4].)
With the advent of 254 nm tunable lasers, a high-performance system was constructed, and the tightly focuses beam allowed the use of higher quantum efficiency small area photodiodes. The small diameter laser beams also meant that the cell walls were not fully bathed in resonance light, plus exposure to Hg was reduced by use of non-resonant detection light. A higher 199Hg density was also possible when using the Faraday effect and lasers, which could be tuned away from the line center.
The cells contained pure CO as a buffer gas, instead of a mixture with N2. Obviously, care was taken to remove oxygen; however, the level of hydrogen contamination is not reported.
These cells were carefully constructed with polished, optically flat end disks and cylinder edges, where a vacuum adhesive was applied. Instead of Varian Torr-Seal epoxy, Space Environment Labs Vacseal silicone-resin-based leak sealant was used. This material has a much lower outgassing rate compared to Torr-Seal, which, together with the close fit between the cylinder and the end-cap disks, minimized the exposure of contaminants that might be generated by the sealant. Furthermore, the cell components were carefully cleaned by refluxing in hydrochloric acid. These cells had stable lifetimes of at least 100 s. There is some indication that the cured silicone resin is slightly permeable to oxygen, which could be a reason for the cell lifetimes to be stabilized. [34]
Four cells were used in this experiment, and later attempts to reproduce these results have not been entirely successful.
I.4 Hydrogen
Since the time that CO was identified as a useful buffer gas, there has been a serious suspicion that cell degradation is due to hydrogen. However, this notion was often met with skepticism, with a counterexample being the Hydrogen Maser: If H atoms were a destructive agent, how could a maser ever operate? In fact, this atomic H is a problem for masers.
One of the first successful maser bulb coatings was Teflon® using DuPont TFK Clear Finish 852-201, while in later work it was applied using DuPont FKP Teflon product code 120. Quote [35]
Although the wall coating procedure described here is quite reliable, the surfaces are not entirely inert, and it is possible that this is due to a contaminant in Teflon. New surfaces are currently being investigated.
In the same year, Berg, one of the coauthors of this work, published a single-author paper stating that excess hydrogen from water-based Teflon coatings and non-UHV vacuum components, rubber o-rings, etc., should be eliminated and had empirical evidence to support that notion.[36]
Later work shows that atomic H actually degrades Teflon coatings, as evidenced by the damage at the atomic beam focus point, on the surface opposite the entrance aperture of the maser bulb.[37]
We encountered similar problems with 199Hg when we used a water-based Teflon coating; coatings made from pure PTFE (polytetrafluoroethylene) in a volatile non-polar carrier give stable and reproducible nuclear spin polarization lifetimes; however, the lifetimes are limited to 15-20 s per cm mfp.
I.4.1 Autocatalytic Hydrogen Production
With the inclusion of CO in the cells, we had hoped that excess hydrogen would form formaldehyde and get absorbed into the wax. As can be seen for the bond energies in Table 1, essentially all materials are susceptible to attack by excited Hg atoms or by atomic H, so the notion of sequestering hydrogen in volatile molecules is not useful.
This leads us to surmise that once hydrogen is present in the system, the rate of generation of more hydrogen actually increases, in the presence of exited-state Hg. That is because each H2 molecule split by an interaction with an Hg will more likely remove another H from a wall coating molecule and form a new H2 molecule. Thus, the degradation process accelerates, and ultimately, the cells, which initially can be rejuvenated by ritualistic treatments involving melting the wax, eventually a useful performance time is shortened to the point that the cells become useless.
Furthermore, atomic H itself is a free radical, and its presence in the cell along with the buffer gas will cause spin relaxation of 199Hg . The lifetime resulting from this will be estimated in the Discussion section.
I.5 Surface States
A previous study, with its carefully acquired large data set, [38], has been very helpful in understanding the dynamics of spin relaxation 199Hg . Here we add to the analysis presented there, first in regard to Eq. (3) of [ 38]. This equation is true if there is a single bound state for an 199Hg atom on the cell surface. However, the number of bound surface states grows with mass,[39], a point that has rarely, if ever, been taken into account in previous similar work (see, e.g., [40]).
The number of states and their energies can be estimated using the WKB approximation, assuming the van der Waals potential (short-range Casimir-Polder force) between an 199Hg atom and a dielectric surface as
$V(z)\approx\cases{}{C}{(d_{0}+z)^{3}}&z\geq 0\\ \infty z<0$ (13)
where $C\approx-6\times 10^{-4}\ e$V nm3 for a heavy polarizable atom near a dielectric surface, with $z$ the distance from the surface in nm and $d_{0}\approx 0.14$ nm representing a sum of radii of the Hg atom and atoms bound in the fused silica surface. The coefficient can be calculated using the formalism in [ 41] with the dielectric function for fused SiO2 in [ 42], Fig. 4, where the calculation was done for Cs. The value for Hg is about 1/6 of that for Cs, due to the shorter wavelength of the allowed electric dipole moment transition of Hg, 180 nm compared to 895 nm and 854 nm (D1 and D2 transitions) for Cs, together with the reduced dielectric response of SiO2 at shorter wavelengths, leading to a reduction in the overlap integral.
The number of bound states is approximately the number of nodes for the least tightly bound WKB solution, which is $n_{max}\approx 72$ for $m=199$, and the two lowest states are approximately -0.18 $e$V and -0.21 $e$V. Here, a plot of $E_{n}$ versus $n$, the number of wavefunction nodes plus one, is shown in Fig. 4 .
Figure 4: Bound state energies as a function of wavefunction nodes plus 1, with a functional approximation to the cumulative distribution function for the states shown by the solid line.
The density of states is a nearly continuum; also plotted is
$F(E_{n})=\frac{N_{max}}{\log(E_{max}/E_{min})}\log(E/E_{max})+1$ (14)
which is a good approximation to a cumulative distribution function of the bound states.
Therefore, Eq. (3) of [38] needs to be modified, and one possibility is to instead take the average sticking time as
$\tau_{s}\propto\sum_{n=1}^{N}\tau_{0n}e^{|E_{an}|/kT}$ (15)
where we allow $\tau_{0}$ to depend on $n$, and $E_{n}$ are the bound state energies. However, since each level is likely to have a different sticking or residence time, a more profitable approach is to calculate the fraction of atoms on the surface by using the partition function.
However, there is no need to consider a sticking time at this point. The binding energies of the levels are occupied according to the Boltzmann distribution, so we can determine the average fraction of the total number of atoms distributed among the bound states. Assume that the atoms move unhindered while in the main cell volume (away from the wall) and at the wall there are bound states with energy $-E_{n}$ (for the following, we take $E_{n}=|E_{n}|$). While in a bound state, an atom can move freely parallel to the surface. Let the volume and area of, for simplicity, a spherical cell be $V$ and $A$, $\beta=1/k_{b}T$, and $p$ be the momentum of an atom. Using simple statistical theory, the probability $\epsilon_{n}$ that an atom spends time in the surface state $n$ is (assuming that this probability is much lower than 1 ),
$\epsilon_{n}$ $=$ $\frac{A\hbar^{-2}\int_{0}^{\sqrt{2mE_{n}}}\exp\left[\left(E_{n}-p^{2}/2m\right)\beta\right]d^{2}p}{V\hbar^{-3}\int_{0}^{\infty}\exp\left(-\beta p^{2}/2m\right)d^{3}p}$ (16)
$\approx$ $\frac{\pi Ae^{\beta E_{n}}\lambda_{T}^{-2}}{V\lambda_{T}^{-3}}=3\pi e^{\beta E_{n}}\frac{\lambda_{T}}{R}$ (17)
where $\lambda_{T}=\hbar(2\pi\beta/m)^{1/2}$ is the thermal wavelength, $\beta=1/k_{b}T$, $m$ is the mass of an atom, and $R$ the radius of the cell, about 1 cm. For Hg at $293\mathrm{~{}K}$, $\lambda_{T}=$ $7\times 10^{-10}\mathrm{~{}cm}$ and using $E_{n}=0.21\ e$V with $R=1\mathrm{~{}cm}$ then $\epsilon=3\times 10^{-5}$.
The total fraction of (temporarily) bound atoms on the surface is given by the sum of $\epsilon_{n}$ for all states. This could be done by summation; however, an approximate analytic formula better illustrates the effects of temperature variation. The bound states as determined by the WKB approximation are shown in Fig. 4, with an estimate of their cumulative probability distribution, the derivative of which gives the density of states. The total fraction of atoms on the wall is therefore
[…]
and the integral is the “exponential integral function," ${\rm Ei}(E_{max}/k_{b}T)-{\rm Ei}(E_{min}/k_{b}T)$, and using Eq. (5.1.10) of 43,
$\epsilon\approx 3\pi\frac{\lambda_{T}}{R}\frac{n_{max}}{\log(E_{max}/E_{min})}\sum_{j=1}^{\infty}\frac{(E_{max}/k_{b}T)^{j}}{j\cdot j!}.$ (19)
[…]
The relaxation rate should depend on the time that an atom is stuck on the surface; however, it is unclear if that is the product of a sticking time and a Boltzmann factor. Given that the time between wall collisions is $\approx v/R$ which is proportional to $\sqrt{T}$, then the relaxation rate should be
$\Gamma_{w}\propto\sqrt{T}\epsilon\propto T\sum_{j=1}^{\infty}\frac{(E_{max}/k_{b}T)^{j}}{j\cdot j!}$ (20)
because $\lambda_{T}\propto 1/\sqrt{T}$. Noting that the derivative of the summation in our original integrand,
$\frac{1}{\Gamma_{p}}\frac{\partial\Gamma_{p}}{\partial T}=\frac{1}{T}-\frac{{e^{E_{max}/k_{b}T}}}{T}\left[{\sum_{j=1}^{\infty}\frac{(E_{max}/k_{b}T)^{j}}{j\cdot j!}}\right]^{-1}.$ (21)
From the graph Fig. 4 of [38], this ratio is about 0.03 at 260 K and 0.02 at 300 K, compared to the formula above numerically determined using $E_{max}=0.21$, with dielectric susceptibility the only specific material properties employed. Of course, correlations times, etc. will depend on temperature, but only weakly. Note that the collision rate times the sticking time $\Gamma_{p}\propto\sqrt{t}e^{\beta E_{0}}$ can be used to fit the data and approximately determines $E_{0}=0.15\ e$V which is close to $0.16\ e$V as obtained in [38] from the data presented in Fig. 4 of that paper.
Note that we did not need to consider a sticking time at all. There is another time scale in the problem, which is the free flight time between wall collision, which for a 1 cm dimension cell is about $10^{-4}$ s. If each traverse across the cell results in a sticking collision, then the time at atom is stuck in the potential wells should be $10^{-4}\epsilon=6\times 10^{-9}$ s to maintain the statistically necessary Boltzmann distribution of atoms. The correlation times in [38] are slightly shorter than this, which we might expect since the atoms move along the surface, the only requirement is that the sticking time be longer than the correlation time; the data in [38] implies the opposite, that the correlation time was less than the sticking time and was left as an open question. If the probability of a wall collision has a probability less than one for an atom to stick, the dwell time becomes longer, as required by microscopic reversibility and the value of $\epsilon$. This implies that the approximately 6 ns time obtained above is strictly a lower limit.
The spatial extent of the surface binding region is only about 0.5 nm, so in $6\times 10^{-9}$ s a 199Hg atom will suffer $10^{4}$ collisions with the wall. Given that there are about 100 surface states, a random walk out of the miasma could very well require $\sim 100^{2}$ steps.
The long wall sticking time for 129Xe as reported in [40] can be explained by there being only a small probability of capture in the wall potential for each wall collision, unlike the case of 199Hg discussed above, where it seems to be of order 50%. Given the chemical reactivity of Hg compared to Xe, this is not surprising, so a sticking probability for Xe between 0.1 and 1 % is needed to explain the results in [40]. The reported binding energy of 0.1 $e$V as determined from the variation of the relaxation rate temperature is probably greater, as in the case of 199Hg where, without the full theory, 0.16 eV was obtained in [38] compared to $\approx 0.21$ eV for the most tightly bound surface obtained from the van der Waals interaction together with the WKB approximation.
Note that $\epsilon\propto 1/R$, whereas the free flight collision time is proportional to $R$, and therefore the overall relaxation rate is independent of $R$, as expected. We can conclude that instead of the characteristic vibration time on the wall, the relevant time scale is $\lambda_{T}/\bar{v}\propto\hbar/k_{b}T$.
The idea of using the vibrational period of an atom near the wall as a sort of sticking time might be applicable for atoms with a bound state spectrum levels below $k_{b}T=0.025\ e$V. This idea was used in [44] from [45], and appears to not obviously apply to 199Hg or 129Xe.
I.5.1 Hydrogen surface states
Again, using the WKB approximation, the surface states of an H atom can be determined. The van der Waals coefficient is 2/3 that of Hg because the principal contributing optical transition for H is at a shorter wavelength. Also, the hard-sphere radius of H is smaller than that of Hg, so we can take $d_{0}\approx 0.12$ nm. Together with its smaller mass, this leads to the result that H has only four bound states at 0.08, 0.03, 0.01, and 0.003 $e$V. Although the energies are smaller than for Hg, the $\epsilon$ is enhanced by a relatively large $\lambda_{T}$ compared to Hg. The net result is $\epsilon=2\times 10^{-6}$, about an order of magnitude smaller than for Hg. Nevertheless, this result implies a microscopically long time for an excited Hg atom to interact with H atoms temporarily stuck to the surface.
We can estimate the effective pressure of H on the surface, assuming that a fraction $62\times 10^{-}$ 6 of H atoms are imprisoned in a region about 0.3 nm thick. If there is 1 mbar H gas in a cell, the number density in the thin region near the cell wall corresponds to a pressure of about 4 bar. Alternatively, with a pressure $P_{0}$ in cell volume, the pressure near the surface (just like the isothermal atmosphere) should be $P_{0}e^{E_{max}/k_{b}T}=P_{s}$, and for $P_{0}=0.001$ bar, with an approximate binding energy of 0.21 $e$ V, it leads to $P_{s}=4.4$ bar and is astonishingly close to the previous estimate of 4 bar.
Relaxation on the wall can be due to an enhanced interaction with H atoms or with the open bonds created when a surface is attacked by atomic H or excited Hg atoms. These open (covalent) bonds carry roughly an electron magnetic moment. The magnetic interaction that leads to relaxation is presented in [38] and can be easily modified for other magnetic moments.
I.6 Relaxation in the Gas phase
Paramagnetic atoms and molecules, which include oxygen and free radicals, can exist in the buffer gas as impurities or the result of 254 nm radiation. Subsequent reactions with excited H can release atomic H (an unpaired spin, so technically a free radical) that exists in the states $F=0$ and $F=1$, the latter that exists as 3/4 of the total H and presents an effective 1/2 of an electron spin due to the addition of the proton nuclear angular momentum.
The relaxation of nuclear spin $I=1/2$ 1S0 atoms by gaseous paramagnetic atoms is theoretically analyzed by Jameson, Jameson, and Hwang [46] for the specific case of O2 and 129Xe, and further developed in [47], for both O2 on 3He and on 129Xe Rewritten here in terms of SI units, and more generally for a colliding molecule (or atom) $X$ with total angular momentum $F$ representing the sum of the internal angular momentum of the constituent nuclei and electron spins of $X$,
$\frac{1}{T_{1}}=\frac{16}{3}F(F+1)~{}\gamma_{I}^{2}~{}\left[\mu_{0}~{}g_{F}~{}\gamma_{e}\right]^{2}\frac{\hbar^{2}}{d^{2}}\left(\frac{\pi\mu}{8k_{B}T}\right)^{1/2}~{}F(U,T)~{}n_{0}\left[X\right]$ (22)
$\gamma_{I}$ and $\gamma_{e}$ are the gyromagnetic ratios for the $I=1/2$ nucleus, and for the bare electron respectively; $g_{F}$ is the Landè $g-$ factor for $X$, and $\mu_{0}=4\times 10^{-7}$. $d$ is the characteristic intermolecular interaction length, $\mu$ is the reduced mass of the colliding species, $k_{B}$ is the Boltzmann constant and $T$ is the absolute temperature (the factor in parentheses is the reciprocal of the mean relative velocity). $n_{0}$ is Loshmidt constant ($n_{0}=2.69\times 10^{25}$/m3 corresponding to the STP number density of an ideal gas), and $[X]$ is the density of the gas $X$ in Amagats. The factor $F(U,T)$ is determined by the interaction potential $U$ between the colliding species, and is unity at high temperature in the hard-core limit. In general, it is determined by integrating solutions of the radial Schr"odinger equation with the potential $U$, together with an average over the Boltzmann momentum distribution. It is noted in [47] that the hyperfine Fermi contact term, proportional to $|\psi(0)|^{4}$, should generally be included in $F(V,T)$, where $\psi(0)$ is the amplitude of the electron wave function of the colliding molecule at the nucleus $I$. The interaction time (correlation time) for gas phase molecules is very short, so effects due to an applied magnetic field become small when the electron precession over the collision time results in a small directional change in the electron(s) spin angle(s); therefore, in all cases under consideration such effects can be ignored for the gas phase.
In [47], the experimental relaxation rate for 3He in O2 is reported to be approximately $0.5$ [s Amagat], which implies $F(U,T)=1.7$. For 129Xe, $F(U,T)\approx 6$ is estimated to be [46] which implies a relaxation rate of 0.7/[s Amagat].
Data in [48] indicates that the relaxation rate of 199Hg in O2 is 800/[s Amagat], which about 1000 times faster than for the inert gases 3He and 129Xe. Using the analysis for inert gases implies that $F(U,T)=1.5\times 10^{4}$ although the two valence electrons of 199Hg likely enhance a hyperfine interaction between the atom and the molecule, the analysis for inert gases is not adequate. Another possibility is that van der Waals molecules can be formed, in which case the needed three-body collisions would make the relaxation rate pressure dependent. This would be interesting to study.
We can try to apply this result to the relaxation of 199Hg by H. Because H has a nuclear spin of 1/2, 3/4 of the atoms will be in the hyperfine level with $F=1$ and $\gamma_{s}$ is reduced by a factor of two compared to O2. Comparing $F(U,T)$ with previous cases of inert gases suggests that $F(U,T)$ scales as the square root of the reduced mass, which we might expect from the Schrödinger equation. Scaling by $\mu$ and $\gamma_{s}^{2}$, and by the fraction of atoms in $F=1$ leads to a relaxation rate of approximately 5/[s Amagat].
The general conclusion is that the pressure of O2 in a room-temperature cell must be below 6 mbar to achieve a spin lifetime of 200 s. In the case of Fig. 1, to achieve a lifetime of 3600 s requires a pressure of 0.3 mbar. This implies that the purity of the gas used to fill the cell to 233 mbar must have been about 99.9%, which is the typical level of commercial gases of non-UHP grade.
A direct measure of the H relaxation rate would require the production of atomic H and a means to measure its pressure. Based on the foregoing, a total H2 pressure less than a few mbar appears as desirable in a cell, particularly given the destructive and autocatalytic behavior of atomic H.
II Materials and Methods
To test the notions that have been discussed so far, the system shown in Fig. 6 was assembled in Fall 1991 at the University of Washington (these data are being reported for the first time). The system consisted of a selection of gases (99 + % purity) that could be introduced into an optical pumping cell contained in a three-layer magnetic shield, together with vacuum pumps and optical pumping components. The cell itself was of volume 750 $\ell$ (we refer to these as “1 liter cells") based on a design used to evaluate the optical pumping of 199Hg in larger volumes than had ever been done before, and was a key R&D component toward the development of the system described in [14].
The cell was coated with a generic pure PTFE lubricating spray in a hydrocarbon carrier. Cells with a coating made using a pure PTFE lubricating spray in a primarily hydro/fluorocarbon carrier (with minor components of isopropyl alcohol and acetone, without water) had better stability than other fluorinated coatings or any water-based Teflon coatings. An example is Elmer’s Slide-All® (which is no longer manufactured); although it was manufactured in a fast-evaporating water-free carrier, it contained a more complex polymer that has hydrogenous functional groups on the ends of PTFE polymer chains, and these groups are subject to attack by excited-state Hg or atomic H. A lubricating spray that uses the same molecule is still available, MicroCare® VDX Dry Lubricant Spray, Aerosol.
Figure 5: Comparison of a non-PTFE lubricating molecule, CAS # 65530-85-0, Poly-tfe omega-hydro-alpha-(methylcyclohexyl) structure to the structure of pure PTFE.
Two coats of a pure polytetrafluoroethylene (PTFE) lubricating spray were applied to the interior surface of the cylinder body, with the edges of the cylinder carefully masked. The end plates were masked to leave a region 1 mm wide from the outer edge of the cylinder uncoated so that the epoxy would have a region to stick between the cylinder wall and the end plate. After coating, the components were baked at 310 C in air and, after cooling, assembled using a minimum of epoxy and avoiding any exposure of epoxy to the inner region of the cell. Using Eq. 1, the mean free path is
$\ell=\frac{4V}{A}=\frac{4\pi R^{2}h}{2\pi R^{2}+2\pi Rh}=\frac{2h}{1+h/R}=6.7\ {\rm cm}$ (23)
where $R$ is the inner radius, and $h$ is the height. These coatings typically gave about 20 s of lifetime per cm mfp, which implies a lifetime of about 130 s for this large cell. The 1 mm fill hole, if complete depolarization occurs for 199Hg atoms entering that region, will lead to a relaxation rate of
$\Gamma_{hole}=\frac{1}{4}\frac{\bar{v}A_{h}}{V}\approx 40\ {\rm s}$ (24)
where $\bar{v}$ is the average Hg velocity and $A_{h}$ is the hole area, as based on the kinetic theory of atomic effusion. The observed cell lifetime was about 136 s, implying that the 199Hg atoms had a good chance of returning to the main volume with their polarization. This is because the capillary inside is coated with PTFE and which, given the smoothness of the surface, indicates that Hg atoms stick briefly to the surface and re-emitted in a random direction.
Figure 6: Manifold and optical system used for relaxation tests. See text for details.
The operation of the system was straightforward. Natural Hg was supplied to the cell from a droplet until the transmitted light was reduced by a factor of 2. Various gases could be admitted and their pressures measured using a Hg manometer or McLeod gauge. A magnetic field of about 10 mG was applied along the light direction to polarize the atoms, and could be suddenly switched perpendicular to the light to observe the modulated light transmission due to the precessing atoms, as measured with a solar-blind photomultiplier. The magnetic shields had a substantial internal magnetic gradient; these shields were constructed from scrap materials and used for testing. The gradient limited the $T_{2}$ lifetime with gases at higher pressures; however, in those cases the longitudinal polarization pump-up and decay could be observed, giving $T_{1}\approx\tau_{0}$, the wall relaxation time.
The pumping was carried out with the full light intensity, which gave a pumping time of 60 s, while the decay was observed with the full light and also with the light intensity reduced by a factor of 10 (using an optical attenuating filter), which allowed extracting the wall decay time $\tau_{w}$.
The main goal of the work was to assess the effect of hydrogen on the wall coating used in [14] which degraded with the application of high voltage, the application of which, we surmised, created atomic H due to inevitable microdischarges on dielectric surfaces. Previous work (see Fig. 1) suggested that oxygen might restore the cell relaxation lifetime. The secondary goal was to test for this possibility and determine how much oxygen could be used before limiting the 199Hg lifetime.
The tests were performed with PTFE instead of deuterated polystyrene (dps) [15] in part because in 1991 the final principal edm cell coating material used in [14] had not yet been decided. Typically, dps provides lifetimes of about 15 s per cm mfp, about as good as PTFE. Resources were not available to continue the 1991 studies with dps.
III Results
The results of a series of measurements are given in Table 2. As can be seen, there were several series of tests. The first was to determine whether the N2, 5% CO buffer gas mix we used in 199Hg edm cells would affect the useful life of the cells or repair damaged cell walls. Tests 1-5 indicate that this gas does not have a direct effect and later damage was not significantly repaired as demonstrated in Tests 12-18.
Test 6 shows that hydrogen gas alone does not damage the wall coating. Tests 7-12 show that Hg in the presence of resonance light immediately degrades the cell, and removal of the gas does not repair the degraded lifetime. Given its stability, the degradation is probably not due to the atomic H being adsorbed on the surface but is more likely due to unterminated bonds on the wall coating. Atomic H has enough energy to do this damage, as shown in Table 1, and reported in [37, 36].
Tests 19-25 show that the addition of oxygen to the cell repairs the damage. We did not verify that resonance radiation is needed to make this repair (neither did we see if hydrogen gas without Hg but resonance light would cause damage).
Tests 15 and 17 indicate that the damage might be local as the buffer gas can tend to hide damage. The 254 nm light entered at exited the cell in a beam of about 2 cm diamter, so it is conceivable that the most significant coating damage was in these areas.
Table 2: Tests conducted with the apparatus shown in Fig. 6. All times have an uncertainty of 8 s unless noted otherwise.
Test No. Condition $\tau_{w}$ [s] Comment/Result
1 only Hg vapor 136 Cell starting lifetime
2 add 1.5 torr N2, 5% CO 133
3 increase to 12 torr " 60 Gradient relaxation
4 increase to 38 torr " 30 "
5 increase to 80 torr " 23 "
6 increase to 112 torr " 16 ", evacuate cell at end of run
5 only Hg vapor 132 cell was evacuated to base pressure
6 10 torr H2 - 10 minutes with no light, evacuate
7 only Hg vapor 132 cell empty other than Hg
8 25 torr H2 30 H2 effect plus Gradient relaxation
9 evacuate to 18 torr 25 "
10 evacuate to 4 torr 15±$\pm 5$ Low signal due to low Hg density
11 evacuate to 0.25 torr 25 Very low signal
12 evacuate 1 hr, add Hg 25 Evident permanent reduction in $\tau_{w}$
13 add 1 torr N2, 5%CO 34 No gradient effect expected
14 increase to 50 torr " no signal
15 increase to 110 torr " 145±$\pm 30$ pump up method so low signal; improvement
with buffer gas implies localized damage
16 evacuate, add Hg 36 partial recovery?
17 add 110 torr N2, 5%CO $\approx 60-100$ similar to 15
18 evacuate, add Hg 36 no apparent further recovery
19 add 20 torr O2 - No observable signal
20 evacuate, add Hg 121 oxygen exposure nearly recovered $\tau_{w}$
21 add 0.65 torr air $\approx 7$ relaxation rate $\approx 1100$/[s Amagat]
22 evacuate, add Hg 134 sat. Hg density, then evacuate to 1/2 abs.
23 evacuated, add Hg 110 evacuation was overnight $2\times 10^{-7}$ torr
24 1 hr irradiation 167 lifetime improving on irradiation
25 no change - stabilized at this value
IV Discussion and Conclusions
Based on the literature on photosensitized Hg reactions, together with our own observations, the results clearly indicate that hydrogen, through atomic H, is nearly certainly responsible for the degradation of 199Hg optically pumped magnetometry cells. Atomic H arises from the reaction of H2 with excited state Hg.
There are two types of cells to consider: closed (permanently sealed) 199Hg EDM cells where the accumulation of impurities cannot be removed without breaking the cell; and open cells used in neutron EDM experiments that employ optically pumped 199Hg as a comagnetometer, which can be evacuated. For the latter, careful attention to the removal of hydrogenous materials, as in the case of the H-maser, which is an open cell system, should help stabilize the 199Hg comagnetometer system. The addition of a small amount of oxygen to the system gave mixed results in [14].
Our 1991 measurement of the relaxation of 199Hg by oxygen gas agrees well with the most recent measurement reported in [48]. Given the salubrious effect of oxygen on cell lifetimes, the inclusion of a small amount will be beneficial; however, it will be consumed to form substances that stick to the wall or mix with wax. In addition, Hg will be consumed, so a source of 199Hg is needed within the cell, or ozone production must be suppressed. Given that of order $10^{13}$ photon per second are incident on a cell and that the photosensitize reaction in some cases has a quantum efficiency near unity, we might expect the same number of atoms or molecules lost per second. This corresponds to $10^{18}$ reaction per day, or 10 mbar of gas produced per day in a 5 $\ell$ cell.
According to Table 1, the only (single) bond that is immune to damage is HF. In addition, the energy available from the Hg bonding to H is available to break other bonds if the vehicle of destruction is HgH.
The addition of CO greatly improved the longevity of cell usefulness, and anecdotal evidence suggests that scrupulous cleanliness and the avoidance of hydrogen gave further improvement, however, the fabrication of cells remains hit-or-miss. A principle effect of CO is to quench the metastable excited state of Hg, which carries enough energy to break bonds.
Hydrogen likely does its damage through an autocatalytic reaction, in that if there is any hydrogen, it will lead to photosensitized production of more hydrogen by breaking the bonds of hydrogenous materials on the cell walls. However, fluorinated waxes appear to not work as well as hydrocarbon waxes, but more studies are needed.
It was hoped that CO might bind to atomic H and eventually form formaldehyde, which would get buried in the wall coating. However, the formyl intermediary is a free radical that probably causes relaxation at a rate similar to that of oxygen, and the hydrogen in both HCO and H2CO are sources for photosensitized hydrogen production.
Identifying the surface states allows for a theoretically consisted picture to explain the data presented in [38]. In all, this was quite remarkable and unexpected.
Similarly, the seemingly anonymously long sticking times for 129Xe as reported in [40] can be explained by such states together with a low probability of sticking per wall collision, which is expected for Xe compared to Hg. Using microscopic reversibility and the calculable value of $\epsilon$ as the fraction of atoms stuck in the wall potential at any given time, if the dwell time can be calculated, perhaps using the Boltzmann transport equation for motion among the myriad of surface states, then the sticking probability is determined. Oddly enough, the reduced interaction of Xe probably means it stays on the surface longer, as the heating interactions that will lead to its release are likely smaller than those for Hg.
Regardless, there is another time scale that is better suited for the case of heavy atoms with many bound states, and that is the thermal wavelength divided by the average velocity, which of course is proportional $\hbar$ divided atom kinetic energy, which instead of the thermal wavelength, is the inverse of the thermal frequency.
An important consequence of the large number of relatively deep surface states (compared to $k_{b}T=0.025$ eV) is in regard to the assumption that a precise volume magnetic field average is provided by 199Hg co-magnetometry. For an extremely precise magnetometry system, as has been proposed, the surface dwell time (during which a substantial spin polarization survives) will need to be more carefully considered, as the dwell time can lead to corrections. For example, localized systematically generated magnetic fields are sensitive to the volume averaging. The gravitation offset of ultracold neutrons in neutron EDM experiments is perhaps a more serious problem; however, all such effects will eventually need consideration.
Based on the foregoing, to improve the durability of closed or open 199Hg optical pumping and detection cells (i.e., any similar cells used for fundamental experiments), some guidelines are as follows:
  1. 1.
    Avoid hydrogen gas either as an impurity in the CO buffer gas (oxygen can be removed, but water as an impurity will produce hydrogen which will pass through the amalgam) or generated by photosensitized reactions. Sealing of filled fused silica cells involves much heat and even UV light is produced, so care is needed in such operations;
  2. 2.
    For open cells used in the neutron EDM experiment that employs a 199Hg comagnetometer, the problem of lifetime degradation appears to be largely solved.[16] This is not surprising as the benzene ring is relatively immune from attack by atomic H [49], and exposure to ozone and oxygen lead to crosslinkages in the polymer.[50]
  3. 3.
    Photosensitized reactions can be suppressed by use of a quenching gas; CO has proven effective but there are better gases;
  4. 4.
    N2O is among the best gas for quenching both 6 3P1 and 6 3P0 excited Hg atoms[51].
  5. 5.
    N2O can be broken by photosensitized reactions with Hg, as is possible with the bond energies presented in Table 1. However, ozone formation, necessary to form HgO, is suppressed due to a scavenging reaction with N2O. Preliminary studies indicated that it might be possible to produce a cell with N2O in a proportion with an inert gas, perhaps a mixture of CO and Xe, that will produce oxygen at a rate sufficient to maintain a long stable life, for example, similar to the data present in Fig. 1, but without the loss of Hg due to the formation of HgO.
  6. 6.
    Fluorinated waxes or silanes with less hydrogen, e.g., Trichloro(3,3,3-trifluoropropyl)silane (CAS 592-09-6). A fully fluorinated silane would be of interest; however, none appear to be commercially available.
  7. 7.
    Deuterated polystyrene[15] remains one of the few options for UCN 199Hg comagnetometer experiments.
Acknowledgments
The author thanks Prof. L.R. Hunter of Amherst College for reporting on the behavior of his cells, Eric Lindahl, the University of Washington glassblower, for technical advice regarding material preparation and for quartzware fabrication, and Jennie Chen for discussions of magnetometry in connection to the LANL EDM project. A comment on the manuscript by Prajwal Mohan Murthy confirming the improvement of $\tau_{w}$ for 199Hg after an oxygen flush is appreciated. I began this work under the direction of Prof. E.N. Fortson in 1982 who was my Ph.D. advisor and Prof. B.R. Heckel, both of the University of Washington, Seattle, WA. Discussion with Prof. J.M. Pendlebury (deceased) of Sussex University led to the recognition of surface states and their role in the behavior of optical pumping cells.
Funding Statement
This research was sponsored by Yale University.
Conflict of Interest
The author declares that there is no conflict of interest.
Institutional Review Board Statement
Not applicable.
Data Availability
All data used in this study are presented in this report or are fully contained in the cited literature.

### v2

@@
-The Chemistry and Physics of 199Hg Nuclear Spin Polarization Relaxation in Quantum Magnetometry Cells
+Surface and chemical effects on ${{}^{1}99}$Hg spin polarization relaxation in optically pumped magnetometers
 Abstract
@@
 As such, a review article is not possible due to the limited specific literature to review.
-Typically, studies of cell atomic polarization relaxation rates (or inverse of the spin lifetime) were done in the context of building an apparatus to measure some fundamental process of greater immediate interest; as typical in such work, the tedious technical details are barely left as a footnote if presented at all.
+Typically, studies of cell wall-induced atomic polarization relaxation times, defined here as $\tau_{w}$ (with wall relaxation rate $\Gamma_{w}=\tau^{-1}_{w}$), were performed in the context of building an apparatus to measure some fundamental process of greater immediate interest; as typical in such work, tedious technical details are barely left as a footnote if presented at all.
 In this research article, optically pumped 199Hg will be put in context with other work in the field.
 In particular, the importance of surface states and the van der Waals interaction of atoms with walls will be discussed together with a theoretical analysis.
-The optical pumping of the ground states of 199Hg and 201Hg was first achieved in the early 1960’s by B.
-Cagnac at the Laboratoire de Spectroscopie Hertzienne de l’ENS (now named the Kastler-Brossel Laboratory).
-[1] The study of the optical pumping of ground state atoms had been expanding through the 1950’s and Hg was among the very atoms first to achieve spin relaxation times of 100’s of seconds.
-(See [2, 3] for a review of the field.) Hg is a 1S0 atom, so the ground-state magnetic moment is due only to the nucleus; for 199 $I=1/2$, and for 201, $I=3/2$.
-Other than application to nuclear magnetic resonance-based gyroscopes [4, 5], optically pumped Hg has found minimal practical applications because alkali atoms, with an unpaired electron spin, provide much greater sensitivities in magnetometer applications.
-In the early 1980’s at the University of Washington, Seattle, Prof.
+The study of the optical pumping of ground state atoms had been expanding through the 1950s and Hg was among the very atoms first to achieve $\tau_{w}$ of 100’s of seconds.
+One of the most important developments in the field is transmission monitoring of the ensemble spin polarization, an idea due to Dehmelt [1] who patented a magnetometer design that is still in use.[2] (See [3, 4] for reviews of the field.)
+The optical pumping of the ground states of 199Hg and 201Hg was first achieved in the early 1960s by B.
+Cagnac at the Laboratoire de Spectroscopie Hertzienne de l’ENS (now named the Kastler-Brossel Laboratory).[5] The theory was fully developed and tested by Cohen-Tannoudji in his Ph.D.
+studies[6, 7], and he has written a general review of the early work at this laboratory.
+[8]
+Hg is a 1S0 atom, so the ground-state magnetic moment, if any, is due only to the nucleus; for the two isotopes with significant natural abundance, only two have non-zero nuclear spin: 199 $I=1/2$, and for 201, $I=3/2$.
+In addition to applications to nuclear magnetic resonance-based gyroscopes [9, 10], optically pumped Hg has found minimal practical applications because alkali atoms, with an unpaired electron spin and hence larger atomic gyromagnetic ratios, can provide much higher sensitivities in magnetometer applications.
+Optically pumped Hg has, however, found application in several fundamental experiments.
+In the early 1980s at the University of Washington, Seattle, Prof.
 E.N.
@@
 These experiments used optical pumping cells that were heated to around 400 C and used the 199Hg nuclear spin as a comagnetometer (a term we invented, using “co" from Latin, meaning “with" or “together") to eliminate frequency changes due to magnetic field fluctuations that would interfere with the fundamental physics measurements via 201Hg.
-However, the main goal of that work was to develop an optical pumping cell in which an electric field could be applied to the processing of 199Hg atoms.
-If the atom has an electric dipole moment in addition to the magnetic one, there would be a shift in the Larmor frequency and would imply the existence of a fundamental time-reversal violating interaction(s) with the 199 Hg atom, the effects of which are expected to be enhanced in high atomic number atoms.
-(See [9] for a review of the field.)
-The goal at the time (1983-5) was to find a cell configuration that would give a long 199Hg spin polarization lifetime and allow the application of a strong electric field to the atoms, without the need to heat the cell to high temperature and also with a buffer gas (both for high voltage stability).
-An earlier experiment in the Fortson group employed spin-exchange polarized 129Xe,[10] for an EDM experiment, and lifetimes near room temperature on the order of a thousand or more seconds were routinely achieved in the presence of a field on the order of a few kV/cm.
-As 199Hg offers a much higher sensitivity to possible fundamental interactions of interest in addition to intrinsically better measurement sensitivity, a vigorous research effort was launched.
-Attempts to find another spin-1/2 atom to serve as a comagnetometer in the same cell as 199Hg were unsuccessful; however, a workable cell was developed with 199Hg alone, although the spin polarization lifetime would degrade with exposure to 254 nm Hg resonance radiation.
-This problem persists to this day in that the construction of stable cells appears as hit-and-miss.
+These experiments were side projects; the main research goal at the time (1983-90) was to develop an optical pumping cell in which an electric field could be generated inside the cell volume and to determine whether the applied electric field leads to a change in the Larmor precession frequency of the nuclear spin of 199Hg atoms.
+If the atom has an electric dipole moment in addition to the magnetic one, there will be a shift in the Larmor frequency and this will imply the existence of a fundamental time-reversal violating interaction(s) with the 199 Hg atom, the effects of which are expected to be enhanced in high atomic number atoms.
+(See [14] for a review of the field, and a more recent review by Chupp et al.
+[15].)
+The essential problem was to find cell coating materials that give a long $\tau_{w}$ for 199Hg and also allow the application of high voltage, without the need to heat the cell to high temperature.
+Inclusion of a relatively inert “buffer gas together with low (room) temperatures is required for high voltage stability.
+An earlier experiment in the Fortson group employed spin-exchange polarized 129Xe,[16] for an EDM experiment, and $\tau_{w}$ near room temperature on the order of a thousand or more seconds were routinely achieved in the presence of a field on the order of a few kV/cm.
+As 199Hg offers a higher sensitivity to possible fundamental interactions of interest in addition to an intrinsically better measurement sensitivity, a vigorous research effort was launched to establish an 199Hg EDM experiment.
+Attempts to find another spin-1/2 atom to serve as a comagnetometer in the same cell as 199Hg were unsuccessful; however, a workable cell was developed with 199Hg alone, although $\tau_{w}$ would degrade with exposure to 254 nm Hg resonance radiation.
+(201Hg has a very short $\tau_{w}$ in room-temperature cells due to its nuclear quadrupole moment, which in addition, anyway, precludes its use in EDM experiments.) This problem persists to this day in that the construction of stable cells remains a hit-and-miss.
+Optically pumped 199Hg has found use in several other fundamental experiments, for example, a tests for axions and their long-range interaction from both spin-polarized and unpolarized material bodies.[17, 18] All of these experiments, including the 199Hg EDM experiments, employ closed cells in that they are filled with appropriate amounts of 199Hg vapor and other gases and permanently sealed off from a preparation manifold.
+In addition, 199Hg has served as a comagnetometer for several neutron EDM experiments, the first application being at the Institut Laue-Langevin.[19, 20] In these experiments, the storage cell for ultra-cold neutrons (UCN) and 199 Hg together needed to be developed, and $\tau_{w}$ stability problems persisted in these experiments.
+A coating that is suitable for both the atomic spin and storage of UCN was found, deuterated polystrerne (DPS) [21], and to date there does not appear to be a fully satisfactory replacement.
+These experiments are “open" in that background gases (including 199Hg) can be pumped out of the cell.
+The techniques were greatly refined by the Paul-Scherrer Institute, [22, 23] leading to the most sensitive neutron EDM measurement performed so far.
+This apparatus was also used to search for axion coupling between a Cu mass and nuclear spins.[24]
+Better wall coatings for 199Hg with longer and stable ensemble nuclear spin polarization relaxation times $\tau_{w}$ are of interest to several ongoing experiments, including the neutron EDM experiments at the Paul-Scherrer Institut [25] and Los Alamos [26], and tests of the equivalence principle by Hunter et al.
+[27] which is the principal motivation for the study presented here.
+The $\tau_{w}$ issues for closed and open cells are related, as will be examined here.
 The purpose of this paper is to briefly review the development of 199Hg optical pumping cells for EDM and other fundamental experiments and to provide an analysis of data that has not been published.
-This data sheds light on the chemical and physical mechanisms of lifetime stability, which, when elucidated, will help identify what is needed to attain reproducibility in cell fabrication.
-Optically pumped 199Hg has found use in several other fundamental experiments, for example, a tests for axions and their long-range interaction from both spin-polarized and unpolarized material bodies.[11, 12] All of these experiments, including the 199Hg EDM experiments, employ closed cells in that they are filled with appropriate amounts of 199Hg and other gases and permanently sealed off from a preparation manifold.
-In addition, 199Hg has served as a comagnetometer for several neutron EDM experiments, the first application being at the Institut Laue-Langevin.[13, 14] In these experiments, the storage cell for ultracold neutrons (UCN) and 199Hg together needed to be developed, and lifetime stability issues persisted in these experiments.
-A coating that is suitable for both the atomic spin and storage of UCN was found [15], and to date, there does not yet appear to be a fully satisfactory replacement.
-These experiments are “open" in that background gases (including 199Hg) can be pumped out of the cell.
-The spin relaxation lifetime issues for the closed and open cells are related, as will be examined here.
-There has been continued interest in neutron EDM and specifically experiments that employ a 199Hg comagnetometer.[16] Although the 199Hg EDM limit[17] is numerically much lower than the neutron EDM (this allows it to be used as a comagnetometer), interpretation of the former results in a sensitivity loss regarding fundamental interactions, so the impacts of both types of experiments on testing the standard model of electroweak interactions are similar.
-New neutron EDM experiments employing 199Hg magnetometry are currently being developed.[18, 19]
+This data sheds light on the chemical and physical mechanisms of the $\tau_{w}$ instability, which, when elucidated, will help identify what is needed to achieve reproducibility in cell fabrication.
+As an aside, although the limit of the 199Hg EDM [28] is numerically much lower than the neutron EDM (this allows it to be used as a comagnetometer), theoretical interpretation of the limits of the 199Hg EDM results in a loss of sensitivity to fundamental interactions, so the impacts of both types of experiments on testing the standard model of electroweak interactions are similar.
+EDM experiments based on optically pumped 199Hg alone have potential for improvement, although the lack of a comagnetometer within the cell volume remains a potential ultimate limitation.
 I.1 Development of 199Hg EDM cells
-Research toward developing a room temperature 199Hg optical pumping cell was begun in 1984, with the first trials using blown fused-silica bulbs of about 18 mm ID.
+The basic optical pumping technique used in the development of 199Hg cells employed circularly polarized resonance radiation and transmission monitoring [1] of the ensemble spin polarization.
+Resonance light from a 204Hg microwave discharge lamp contains a single 254 nm line (corresponding to the $6^{1}S_{0}$ to $6^{3}P_{1}$ transition) because there is no hyperfine structure for this isotope.[5, 7] The single spectral line overlaps the $F=1/2$ ground state to $F=1/2$ excited state levels of 199Hg.
+If a cell containing atoms is placed in a few mG magnetic field oriented along the light propagation direction, the ground state atoms will be excited only from the $m_{F}=-1/2$ if the light has left-handed circular polarization (which carries $+1$ unit of angular momentum).
+The decay back to the ground state can go to either level; however, atoms accumulate in the $m_{F}=1/2$ level because these never go through an optical pumping cycle.
+Eventually, the atomic nuclear spin polarization reaches equilibrium, and the transparency of the cell increases.
+The equilibrium spin polarization is, when the only spin polarization relaxation is due to wall interactions, is
+$P=\frac{\Gamma_{p}}{\Gamma_{p}+\Gamma_{w}}$ (1)
+where $\Gamma_{p}$ is the optical pumping rate or the inverse of the pumping cycle time, $1/\tau_{p}$, and $\Gamma_{w}=1/\tau_{w}$ is the wall relaxation rate.
+For most of the subsequent discussion, it will be assumed that wall relaxation is dominant over magnetic gradient effects or interactions with other atoms or molecules in the gas phase.
+In the weak field limit ($\ll 100$ G), in the absence of magnetic gradients, the longitudinal and transverse (coherence) relaxation times are equal ($T_{1}$ and $T_{2}$ in nuclear magnetic resonance terminology).
+(See [3] for a comprehensive review of ground-state optical pumping, in which Fig.
+1 is a pictorial of this specific case.)
+If the magnetic field is suddenly switched to be perpendicular to the light (in time much faster than the inverse Larmor frequency), $P$ becomes time dependent, and the transmitted light is modulated at the Larmor frequency, which is about 7 Hz for a 10 mG magnetic field.
+The modulation level will exponentially decay with a total rate given by the sum of $\Gamma_{p}$ and $\Gamma_{w}$, so to determine $\tau_{w}$, the total rate must be extrapolated to $\Gamma_{p}=0$.
+For discharge lamps, the non-resonant background light was typically only 20% of the emitted light, with the dominant noise source being the shot noise in the detected photons.
+Recent experiments have employed frequency-quadrupled semiconductor lasers, which have reduced non-resonant light and higher-efficiency detectors.
+Furthermore, such lasers are tunable, which allows for a higher density of 199Hg because the laser can be tuned away from the absorption line center.
+Research toward developing a room temperature 199Hg optical pumping cell was begun in 1984, with the first trials using fused-silica blown bulbs of about 18 mm ID.
 The kinetic mean free path (mfp) between wall collisions is given by
 $\ell=\frac{4V}{A}=\frac{4}{3}R$ (2)
-where $V$ and $A$ are the internal volume and surface area, and $R$ is the sphere radius.
+where $V$ and $A$ are the internal volume and surface area, and $R$ is the radius of the sphere.
 These cells also had fill stems about 20 mm long where they were sealed from the vacuum system after being filled with natural Hg 0 C vapor pressure.
-The first cell had the inner surface coated with paraffin (dotriacontane) and had a spin polarization lifetime of about 40 s.
-For a larger cell as envisioned for EDM work, with a factor of 2 greater mfp, assuming homogeneous surface relaxation, the lifetime would be 80 s, and good enough for an EDM experiment.
-Under illumination with the 254 nm pumping light, the cell lifetime deteriorated.
-It could be temporarily restored by remelting the wax; overall, ominous effects that we did not fully appreciate at the time.
-The next bath of spherical cells had siliconizing agents applied to the inner surface, both Surfasil® and Aquasil® were tested, both with and without wax, and one Aquasil® cell was filled with 40 torr N2 buffer gas (Linde® spectroscopic grade, of unknown history).
+The first cell had the inner surface coated with paraffin (dotriacontane) “wax" and had a $\tau_{w}$ of about 40 s.
+For a larger cell, as envisioned for EDM work, with a factor of 2 greater mfp, assuming homogeneous surface relaxation, $\tau_{w}$ would be 80 s and good enough for an EDM experiment.
+Under illumination with the 254 nm pumping light, the cell’s $\tau_{w}$ deteriorated.
+It could be temporarily restored by remelting the wax; this potentially ominous effect was not fully appreciated at the time.
+The next batch of spherical cells had siliconizing agents applied to the inner surface, both Surfasil® and Aquasil® were tested, both with and without wax, and an Aquasil® cell was filled with 40 torr N2 buffer gas (Linde® spectroscopic grade).
 The wax was introduced only after the cells without wax were removed from the vacuum system.
 It was observed that the wax did not wet the siliconized surfaces but accumulated in individual droplets, so the wax was driven into the cell stem, where it formed a sort of crystalline plug.
-All of these cells had a lifetime of only about 15 sec.
-This was thought to be due to the damage to the coating that occurred with the cells being sealed off, the poor character of the wax plug, or magnetic field gradients in the case of the buffer gas cell (the magnetic shield was very crude), which had a 5 s.
-lifetime.
+All of these cells had a $\tau_{w}$ of only about 15 s.
+This was thought to be due to the damage to the coating that occurred with the cells being sealed off, the poor character of the wax plug, or magnetic field gradients in the case of the buffer gas cell (the magnetic shield was very crude), which had a 5 s $\tau_{w}$.
 The Aquasil® coated cell (10% dilution in acetone, followed by two rinses with acetone and drying by blowing air into the bulb) had the best long-term performance.
-Warming the cell to 120 $\circ$ C increased useful life to 60 s, with a modest permanent improvement at room temperature.
-Emboldened by these results, we made cells similar to those used in [10] but with Hg and only N2 as a buffer gas that is required for high-voltage stability.
-These cells had a 25 mm OD, 23 mm ID, 10 cm tall fused silica cylinder with a fill stem that enters radially into the side of the cylinder, and capped by fused silica disks of approximately 38mm diameter, glued on with Varian Torr-Seal® epoxy.
-For Hg, a variety of electrode materials were tried; however, exactly the same electrode coating material (tin oxide) coated with wax with Aquasil® on the cylinder body gave excellent performance.
-In Fig.
-1, a free-precession signal is shown, obtained by optical pumping with resonance light from a 204Hg microwave discharge lame; there is no hyperfine structure for this isotope, and the single spectral line overlaps the $F=1/2$ to $F=1/2$ level for 199Hg, and the $F=3/2$ to $F=5/2$ for 201Hg (for which no optical pumping signal was ever observed in our EDM cells, as expected).
-This plot was obtained by optically pumping the system with circularly polarized resonance light propagating along the direction of a weak magnetic field (10 mG), which was suddenly switched off with the simultaneous sudden application of a second field (10 mG) perpendicular to the light.
-The angular momentum states for the transition in 199Hg are optimal for optical pumping because the sample becomes transparent (for pure 199Hg) when fully polarized, and the modulation on the light during free precession is 100%, however that modulation is reduced to about 15% with a natural isotopic mixture of Hg.(See Fig.
-1 of [2] for a pictorial of this specific case.)
-Remarkably, the relaxation time in the dark was about 3600 s.
+Warming the cell to 120 C increased useful life to 60 s, with a modest permanent improvement at room temperature.
+Emboldened by these results, we made cells similar to those used in the 129Xe experiment [16] but with Hg and only N2 as a buffer gas required for high-voltage stability.
+These cells had a fused silica cylinder of 25 mm OD, 23 mm ID, 10 mm high with a fill stem that enters radially onto the side of the cylinder, and the cylinder was capped with fused silica disks of approximately 38 mm diameter, glued with Varian Torr-Seal® epoxy.
+To provide a means of high voltage (HV) application within the cell to establish an electric field, a variety of electrode materials were tried; however, exactly the same electrode coating material (tin oxide) as used in the 129Xe experiment, coated with wax, and with Aquasil® on the cylinder body gave excellent performance.
+The data shown in Fig.
+1 was obtained by optically pumping the system with circularly polarized resonance light propagating along the direction of a weak magnetic field (10 mG), which was suddenly switched off with the simultaneous sudden application of a second field (10 mG) perpendicular to the light, as described earlier.
+The angular momentum states for the transition in 199Hg are optimal for optical pumping because the sample becomes transparent (for pure 199Hg) when fully polarized, and the modulation on the light during free precession can be nearly 100% for a fully polarized ensemble; however, the modulation is reduced to about 15% with a natural isotopic mixture of Hg, as was used throughout the cell development period and for the first 199Hg EDM experiment.
+The actual Larmor frequency was about 10 Hz, which was mixed with the signal from a constant oscillator to bring the precession frequency down to where it could be measured with a chart recorder.
+At the minimum controllable light level that we could achieve with our system, $\tau_{w}$ was 300 s, and met the 100 s goal for cell development (because $1/f$ noise magnetic field noise required a relatively fast electric field reversal rate, so a longer time was unnecessary).
+A quick test in the dark indicated a lifetime in excess of 1000 s, which is consistent with other ${}^{1}S_{0}$ nuclear-spin 1/2 atoms, particularly 129Xe.
 We therefore thought that we were done with cell development, as cells made in this manner proved to be stable under the application of high voltage.
-Figure 1: 1st successful EDM cell, probe light limited lifetime to 200 sec, dark lifetime was 3600 s at the time of the measurement.
+Figure 1: 1st successful EDM cell, probe light limited the spin polarization relaxation time to 200 s.
 Unfortunately, our hopes were dashed when we discovered that the cells were unstable against the loss of Hg.
@@
 The phase shift between the atomic precession and the oscillating field is proportional to the degree of detuning between the Larmor and oscillating field frequencies and can be measured with a phase-sensitive detector.
-Two such cells were operated together with electric fields directed oppositely, with the sum of the frequency difference used to stabilize the applied field and a correction signal to stabilize the gradient field and which provided a signal to detect an EDM because the field in the cells was directed oppositely.[20]
-The in-phase transmitted light signal provides a measure of the sample atomic polarization, times the atomic density.
+Two such cells were operated together with electric fields directed oppositely, with the sum of the frequency difference used to stabilize the average applied field, and the difference used as a correction signal to stabilize the gradient field between the cells.
+The gradient correction signal was sensitive to an EDM because the electric field in the cells was directed oppositely.[29]
+The in-phase oscillating transmitted light signal provides a measure of the sample atomic polarization, times the atomic density.
 Shown in Fig.
-2 is one of the data runs where the loss of Hg one of the cells is evident as the average transmitted light intensity, measured by a solar blind photomultiplier, steadily increases, representing a loss of Hg atoms (the discharge lamps used in this work were remarkably stable, with few percent at most change in 24 hours).
-At the same time, the in-phase signal substantially increases, and this is only possible if the wall lifetime increases, which increases the atomic polarization for the approximately constant optical pumping rate.
-Figure 2: $V_{DC}$ is a measure of the transmitted light intensity (multiplied by 10 to offset the plot from $V_{in}$), and its increase indicates a loss of Hg atoms.
-The atomic oscillator in phase signal continues to increase, which is only possible if the cell relaxation rate decreases (e.g.
-lifetime increases), and then disappears when the Hg density falls toward zero.
-Data manually digitized from plots in [7], Fig.
+2 is one of the data runs in which the loss of Hg in the cells is evident at the end of the run when the oscillator signal $V_{in}$ fades.
+Figure 2: $V_{DC}$ is a measure of the transmitted light intensity (multiplied here by 10 to offset the plot from $V_{in}$), and its increase indicates a general upward drift of the light intensity, with a rapid loss of Hg atoms near the end time, as further evidence by the loss of the oscillator in-phase signal $V_{in}$.
+This data was manually digitized from plots in [12], Fig.
 8.2, cell 2.
-For the case of low absorption, the transmitted light is, by expanding $e^{-x}\approx 1-x$,
-$I(t)=I_{0}\left[1-\rho(t)\sigma\ell\big{(}1-\epsilon P(t)\cos\omega t\big{)}\right]$ (2)
-where $I_{0}$ is the resonant light intensity, $\rho(t)$ is the time-dependent Hg density, $\sigma$ is the light absorption cross section, $\ell$ is the effective path length for light through the cell, $\epsilon$ is the fraction of light resonant with 199Hg , $P(t)$ is the average magnitude of the atomic polarization, and $\omega$ is the Larmor frequency.
-Therefore the average $DC$ light intensity, represented by $V_{DC}(t)$, is
-$V_{DC}(t)\propto I_{0}\big{(}1-\rho(t)\sigma\ell\big{)}$ (3)
-and the in-phase signal is
-$V_{in}\propto I_{0}\rho(t)\sigma\ell P(t).$ (4)
-The fractional change in the amount of light absorbed by the atoms, relative to the initial, is
-$\eta(t)=\frac{V_{DC}(\infty)-V_{DC}(t)}{V_{DC}(\infty)-V_{DC}(0)}$ (5)
-$V_{dc}(\infty)\approx V_{dc}(t_{f})+30$ (6)
-in the voltage units of Fig.
-2, and where $t_{f}=18000$ s is the end of the measurement time however the absorption has not yet reached zero.
-A rough extrapolation gives a small increase in the absorption.
-It is the absorbed light that produces $V_{in}$, so
-$\frac{V_{in}(t)}{\eta(t)}\propto P(t).$ (7)
-From the basic principles of the atomic oscillator,
-$P(t)=\frac{\Gamma_{p}}{\Gamma_{p}+\Gamma_{w}+\omega_{1}}$ (8)
-where $\Gamma_{p}$ is the optical pumping rate, $\Gamma_{w}=1/\tau_{w}$ is the spin relaxation rate and $\omega_{1}$ is the effective precession rate around the oscillating field.
-Because the in-phase atomic oscillator signal $V_{in}(t)$ is proportional to the absorbed light $\eta(t)$, it follows that
-$\frac{\eta(t)}{V_{in}(t)}\propto(P(t))^{-1}=\frac{\Gamma_{p}(t)+\Gamma_{w}(t)+\omega_{1}}{\Gamma_{p}(t)}.$ (9)
-The pumping rate is given by the average DC light intensity in the cell, which is
-$\Gamma_{p}\propto[1-\alpha\eta(t)]$ (10)
-where $\alpha=0.58$ is the initial light absorption as determined by $V_{dc}(0)$ and $V_{dc}(\infty)$, noting that typically 20% of the light is nonresonant for the microwave discharge lamps we employed.
-Therefore,
-$\frac{[1-\alpha\eta(t)]\eta(t)}{V_{in}(t)}\propto\Gamma_{p}(t)+\Gamma_{w}(t)+\omega_{1}=\Gamma_{tot}(t).$ (11)
-We can take the ratio $\Gamma_{tot}(t)/\Gamma_{tot}(0)$ to determine how the total relaxation lifetime evolves, which is shown in Fig.
-3 to decrease.
-This can be compared to the pumping rate $\Gamma_{p}(t)/\Gamma_{p}(0)$, which increases over the same time period, because the light intensity in the cell increases as Hg is lost.
-The other term in $\Gamma_{tot}(t)$ is $\omega_{1}$ which is constant.
-In typical operation of the atomic oscillators, the light intensity and $\omega_{1}$ were kept constant between data runs. Normally we tried to achieve
-$\omega_{1}=\Gamma_{p}\approx\Gamma_{w}/2.$ (12)
-This relationship can only be satisfied for the data in Fig.
-3 if $\Gamma_{w}(0)>5\Gamma_{p}$, a factor 2.5 times larger than the target value.
-However, this choice provided a reasonable balance over the run, plus stronger pumping tended to degrade the cells faster.
-This shows that, without a doubt, during this data run, $\Gamma_{tot}(t)$ decreased with time, so $\tau_{w}$ must have increased as Hg was consumed.
-From this graph, we estimate that $\tau_{w}$ increased by at least a factor of three over the duration of the experimental run (data near the end of the run are unreliable due to the failure of the atomic oscillator as the signal diminished) to within the approximations and the lack of experimental specifics for work done nearly 40 years ago.
-However, this is the first clear evidence that 254 nm resonance radiation irradiation can be beneficial in one way, and perhaps the only direct evidence to date.
-In principle, a source of Hg in the cell could be along with a small amount of oxygen; however, the optical density would be too high for any practical system at the time.
-Figure 3: By use of the cell light transmission, it is possible to deduce the the pump rate $\Gamma_{p}$, and the in-phase signal to deduce $\Gamma_{p}+\Gamma_{w}+\omega_{1}$.
-As the last term is constant, and their sum is decreasing while $\Gamma_{p}$ is increasing, is clear that the wall relaxation rate is decreasing.
-The loss of Hg atoms eventually lead to a failure of the atomic oscillator for this cell, near the final time of the plotted data.
-In obtaining the first 199Hg EDM limit, new cells were manufactured for each short data run and the result of that was published [7, 20]. .
+This data has some interesting features that can be understood by considering the degree of polarization of the ensemble,
+$P=\frac{\Gamma_{p}}{\Gamma_{p}+\Gamma_{w}+\omega_{1}}$ (3)
+where $\omega_{1}$ is the oscillating field amplitude.
+Normally, the light intensity in our experimental work was very stable, so this data run is slightly anomalous; however, given the haste with which we had to set up and test the system, it is understandable.
+At the start of the experiment, the light intensity and $\omega_{1}$ were adjusted such that $\Gamma_{p}\approx\omega_{1}$ and $\Gamma_{p}+\omega_{1}\approx 50\ {\rm s}^{-1}$, with approximately 30% precision.
+Noting that the light intensity is proportional to $\Gamma_{p}$, and that $\Gamma_{w}=\tau_{w}^{-1}\ll\Gamma_{p}$, the in-phase signal should scale with the light intensity as its product with $P$:
+$V_{in}\propto\frac{\Gamma_{p}^{2}}{\Gamma_{p}+\omega_{1}}\approx\frac{\Gamma_{p}}{2}.$ (4)
+This scaling works up to about 13000 seconds, at which time $V_{DC}$ begins to increase rapidly, while $V_{in}$ begins to fall.
+This marks the onset of the loss of Hg due to reactions with atomic oxygen, or more accurately, ozone O3 which takes some time to build up.
+This process will be described in some detail later in this article; is a two-step process by which Hg reacts with O2 to form O3 and atomic O.
+O3 reacts irreversibly with Hg to form HgO, resulting in an increase in transmitted light.
+The initial O3 concentration is zero, so the loss of Hg does not begin until it is formed, with the loss rate increasing over irradiation time.
+The differential equation governing this has a loss rate that increases linearly with time (neglecting the loss of Hg so this is approximate, for the early time development),
+$\frac{d\rho}{dt}=-\alpha t\rho$ (5)
+where $\rho$ is the Hg density, and $\alpha$ is a constant rate parameter that is proportional to the resonance light intensity.
+A full model with both varying Hg density and light intensity gives a good representation of the observed time dependence.
+In the case of constant intensity, the solution is $\rho(t)=\rho(0)e^{-\alpha t^{2}}$ which explains the sudden drop in $V_{in}$ at the end of the run.
+If a cell had been irradiated, the density would continue to change as $\rho(t)=\rho(t_{0})e^{-\alpha t_{0}t}$, assuming the ozone concentration is constant.
+It can be further noted that the change in $V_{DC}$ when the Hg disappears, accounting for the upward drift in light intensity, implies that the Hg light absorption at the beginning of the run was only about 15%.
+This is most likely due to an earlier loss of Hg from the same photosynthesized reaction.
+Even if a cell is briefly illuminated, there will be some ozone formation, and although at low concentrations the rate of Hg loss is slow, nonetheless the reactions will continue until the ozone is absorbed or reacted with the materials within the cell, as described above.
+Note that these cells were operated in the strong pumping regime, so it is not possible to deduce if there was a change in $\tau_{w}$ based on these data.
+The loss of Hg limited the duration of the EDM measurements for each cell pair, which for the case discussed was about 5 h.
+In obtaining the first 199Hg EDM limit, we opted to manufacture new cells (in batches of eight) that would be used in pairs until their Hg density decreased to a point where the oscillators failed.
+The net result of this imperfect but successful study was published.[12, 29]
+In principle, a source of Hg could have been incorporated into cells and thus keep the Hg density constant; however, the optical density would have been too high for any practical system at the time, as the proper natural Hg density roughly corresponds to the $0^{\circ}$ vapor pressure.
+(199Hg accounts for only 15% of the total absorption of light produced by 204Hg; later cells employed enriched 199Hg .) At the time, no tunable sources of resonance light were available.
+This allows probing away from the line center, as is done currently using quadrupled semiconductor lasers.
 Table 1: Excitation and bond energies of interest. Most are available in chemistry textbooks (e.g., [30]); citations are provided for the rarely encountered bonds.
@@
 or Bond (enthalpy) $[e{\rm V}]$
-3P1 Hg 4.89 allowed transition ($f\approx 1$)
+3P1 Hg 4.89 allowed transition (oscillator strength $f\approx 1$)
 3P0 Hg 4.55 metastable, populated by N2 collisions with 3P1 Hg
@@
 C-H 4.26
+(H5C6)-H 4.46 extra binding energy compared to C-H due to resonance
+stabilization of the phenyl ring
 F-H 5.80
@@
 There is no isotopic enrichment in the HgO formed in this process, which led to the full elucidation of this process [37] that has been known for close to 100 years.[38] It can be understood that such a series of reactions is possible from the bond energies given in Table 1.
-Armed with this knowledge, we took steps to remove oxygen from the nitrogen buffer gas by bubbling the gas through liquid amalgams of Mg-Hg or Al-Hg followed by a liquid nitrogen trap to remove Hg vapor.[30] After such purification, the Hg density became constant; however, the cell lifetimes would rapidly degrade with resonance light irradiation.
-A further study of photosensitized reactions indicated that metastable 6 3P0 atoms, which are readily formed with collisions between 6 3P1 atoms and nitrogen molecules, due to resonance with the rotation states of N2.
+Armed with this knowledge, we took steps to remove oxygen from the nitrogen buffer gas by bubbling the gas through liquid amalgams of Mg-Hg or Al-Hg followed by a liquid nitrogen trap to remove Hg vapor.[39] After such purification, the Hg density became constant; however, the cell’s $\tau_{w}$ rapidly degraded with resonance light irradiation.
+A further literature study of photosensitized reactions indicated that metastable 6 3P0 Hg atoms can be readily formed with collisions between 6 3P1 Hg atoms and nitrogen molecules, due to resonance with the rotational states of N2.
 The 6 3P0 excited atoms still carry considerable energy, see Table 1, and can damage the wall coatings.
-The metastable lifetime has been estimated as seconds, so there is ample time for diffusion to the cell wall.
+The metastable atomic state lifetime has been estimated as a few seconds, so there is ample time for diffusion to the cell wall.
 The destructive nature of 6 3P1 and 6 3P0 atoms has been known for nearly 100 years (this was known before the possibility of producing ozone [40]).
@@
 Indeed, even bare fused silica is prone to damage by these excited atoms.
-In our earlier work with testing spatial isotropy, we discovered that fused silica optical pumping cells that were irradiated while at room temperature would suffer irreversible damage as manifested by a shorted spin lifetime.
-The damaging effects have been known since 1930 [32] but apparently not by us.
-There was a vague notion that hydrogen was responsible for the degradation of the spin lifetime; however, a model was never developed, and the notion tended to be dismissed because of the successful operation of hydrogen masers (more on this later).
+In our earlier work with testing spatial isotropy, we discovered that fused silica optical pumping cells that were irradiated while at room temperature would suffer irreversible damage as manifested by a shorted $\tau_{w}$.
+The damaging effects on silica have been known since 1930 [41] but apparently not by us.
+There was a vague notion that hydrogen was responsible for the degradation of the $\tau_{w}$; however, a model was never developed and the notion tended to be dismissed due to the successful operation of hydrogen masers (more on this later).
 The best option to quench 6 3P1 (that is, perturb the excited state causing a radiative transition or energy transfer) directly to the ground state and also to quench 6 3P0 (which is not quenched by nitrogen) is carbon monoxide, CO.
-Cells were made with 5% CO and 95%N2 and the cell life expectancy improved.
-UW 199Hg EDM experiments over the next decade employed such cells.
+Cells were made with 5% CO and 95% N2 and the useful life expectancy of cells under UV irratiation improved.
+The University of Washington 199Hg EDM experiments over the next decade employed such cells.
 I.3 Further Advances
-The latest 199Hg EDM experiment is reported in [17], and incorporated a number of improvements.
-First, it used a pulsed pump and probe method that gave a factor of nearly 2 in the average atomic polarization, as suggested in [33].
-This was demonstrated together with spin precession detected by the Faraday rotation of the off-resonant light provided by an 200Hg discharge lamp.
+The latest 199Hg EDM experiment is reported in [28], and incorporated several improvements.
+First, it used a pulsed pump and probe method that gave a factor of nearly 2 in the average atomic polarization.
+The spin precession was detected by use of the Faraday effect (linear polarization rotation) of second light beam that was not resonant with any 199Hg 254 nm transitions.
+Thus, there was little or no optical pumping relaxation due to the detection light, and the pump light was blocked during a measurement.
 (It should be noted that a similar system is described in [9].)
-With the advent of 254 nm tunable lasers, a high-performance system was constructed, and the tightly focuses beam allowed the use of higher quantum efficiency small area photodiodes.
-The small diameter laser beams also meant that the cell walls were not fully bathed in resonance light, plus exposure to Hg was reduced by use of non-resonant detection light.
+This high performance system was made possible by advent of 254 nm tunable lasers, and the tightly focused beams allowed the use of higher quantum efficiency small area photodiodes.
+The small diameter laser beams also meant that the cell walls were not fully bathed in resonance light, and exposure to Hg was reduced by the use of non-resonant detection light.
 A higher 199Hg density was also possible when using the Faraday effect and lasers, which could be tuned away from the line center.
-The cells contained pure CO as a buffer gas, instead of a mixture with N2. Obviously, care was taken to remove oxygen; however, the level of hydrogen contamination is not reported.
+The cells contained pure CO as a buffer gas, instead of a mixture with N2.
+Obviously, care was taken to remove oxygen; however, the level of hydrogen contamination is not reported.
 These cells were carefully constructed with polished, optically flat end disks and cylinder edges, where a vacuum adhesive was applied.
@@
 This material has a much lower outgassing rate compared to Torr-Seal, which, together with the close fit between the cylinder and the end-cap disks, minimized the exposure of contaminants that might be generated by the sealant.
-Furthermore, the cell components were carefully cleaned by refluxing in hydrochloric acid.
-These cells had stable lifetimes of at least 100 s.
-There is some indication that the cured silicone resin is slightly permeable to oxygen, which could be a reason for the cell lifetimes to be stabilized.
+The cell components were carefully cleaned by refluxing in hydrochloric acid.
+These cells had a stable $\tau_{w}$ of at least 100 s.
+This is notably lower than the 300 s light-limited spin polarization relaxation time obtained in the first EDM cells which had a small amount of O2 gas.
+Although the presence of O2 resulted in a loss in Hg (gas), it apparently improved the spin relaxation properties of the surface.
+It is well known that some cured silicone resins are slightly permeable to oxygen, [42] which could be a reason for the stabilization of the cell spin polarization relaxation time.
 [43]
@@
 I.4 Hydrogen
-Since the time that CO was identified as a useful buffer gas, there has been a serious suspicion that cell degradation is due to hydrogen.
+Since CO was identified as a useful buffer gas, there has been a serious suspicion that the cell degradation is due to hydrogen.
 However, this notion was often met with skepticism, with a counterexample being the Hydrogen Maser: If H atoms were a destructive agent, how could a maser ever operate?
@@
 New surfaces are currently being investigated.
-In the same year, Berg, one of the coauthors of this work, published a single-author paper stating that excess hydrogen from water-based Teflon coatings and non-UHV vacuum components, rubber o-rings, etc., should be eliminated and had empirical evidence to support that notion.[36]
-Later work shows that atomic H actually degrades Teflon coatings, as evidenced by the damage at the atomic beam focus point, on the surface opposite the entrance aperture of the maser bulb.[37]
-We encountered similar problems with 199Hg when we used a water-based Teflon coating; coatings made from pure PTFE (polytetrafluoroethylene) in a volatile non-polar carrier give stable and reproducible nuclear spin polarization lifetimes; however, the lifetimes are limited to 15-20 s per cm mfp.
+In the same year, Berg, one of the coauthors of this work, published a single-author follow-up paper stating that excess hydrogen from water-based Teflon coatings and non-UHV compatible vacuum components (rubber o rings, etc.) should be eliminated and had empirical evidence to support that notion.[45]
+Later work shows that atomic H actually visibly degrades Teflon coatings, as evidenced by the damage at the atomic beam focus point, on the surface opposite the entrance aperture of the maser bulb.[46]
+We encountered similar problems with 199Hg when we used a water-based Teflon coating; coatings made from pure PTFE (polytetrafluoroethylene) in a volatile non-polar carrier give stable and reproducible $\tau_{w}$; however, the $\tau_{w}$ are limited to 15-20 s per cm mfp.
 I.4.1 Autocatalytic Hydrogen Production
@@
 This leads us to surmise that once hydrogen is present in the system, the rate of generation of more hydrogen actually increases, in the presence of exited-state Hg.
-That is because each H2 molecule split by an interaction with an Hg will more likely remove another H from a wall coating molecule and form a new H2 molecule.
-Thus, the degradation process accelerates, and ultimately, the cells, which initially can be rejuvenated by ritualistic treatments involving melting the wax, eventually a useful performance time is shortened to the point that the cells become useless.
-Furthermore, atomic H itself is a free radical, and its presence in the cell along with the buffer gas will cause spin relaxation of 199Hg .
-The lifetime resulting from this will be estimated in the Discussion section.
+This is because the two H atoms from the splitting of each molecule split by an interaction with an Hg will more likely combine with and remove H atoms from a wall coating molecule and form new H2 molecules, rather than finding and recombining with their former partners.
+Thus, the degradation process accelerates and, ultimately, cells that initially could be rejuvenated by ritualistic treatments that involve melting wax, eventually had $\tau_{w}$ degraded to the point that the cells become useless.
+Furthermore, atomic H itself is a free radical, and its presence in the cell along with the buffer gas can cause the relaxation of spin polarization of 199Hg , which adds to that due to the walls.
+The spin polarization relaxation time resulting from this is estimated in the discussion section.
 I.5 Surface States
-A previous study, with its carefully acquired large data set, [38], has been very helpful in understanding the dynamics of spin relaxation 199Hg .
-Here we add to the analysis presented there, first in regard to Eq.
+A previous study, with its carefully acquired large data set, [47], has been very helpful in understanding the dynamics of spin relaxation of 199Hg .
+Here we add to the analysis presented there and recapitulate and expand the analysis presented in [48].
+First, consider Eq.
 (3) of  [47].
 This equation is true if there is a single bound state for an 199Hg atom on the cell surface.
-However, the number of bound surface states grows with mass,[39], a point that has rarely, if ever, been taken into account in previous similar work (see, e.g., [40]).
-The number of states and their energies can be estimated using the WKB approximation, assuming the van der Waals potential (short-range Casimir-Polder force) between an 199Hg atom and a dielectric surface as
+However, the number of bound surface states grows with the mass of the bound atom,[49], a point that has rarely, if ever, been taken into account in previous similar work (see, e.g., [50]).
+The number of states and their energies can be estimated using the Wentzel-Kramers-Brillouin (WKB) approximation, which is an approximate method for determining the bound state energies and the wavefunctions of a massive particle bound in a spatially varying potential[51].
+The van der Waals potential (short-range Casimir-Polder force) between an 199Hg atom and a dielectric surface is given by
 $V(z)\approx\cases{}{C}{(d_{0}+z)^{3}}&z\geq 0\\ \infty z<0$ (6)
-where $C\approx-6\times 10^{-4}\ e$V nm3 for a heavy polarizable atom near a dielectric surface, with $z$ the distance from the surface in nm and $d_{0}\approx 0.14$ nm representing a sum of radii of the Hg atom and atoms bound in the fused silica surface.
+where $C\approx-6\times 10^{-4}\ e$V nm3 for a heavy polarizable atom near a dielectric surface, with $z$ the distance from the surface in nm and $d_{0}\approx 0.14$ nm representing the sum of radii of the Hg atom and atoms bound in the fused silica surface.
 The coefficient can be calculated using the formalism in  [52] with the dielectric function for fused SiO2 in  [53], Fig.
 4, where the calculation was done for Cs.
-The value for Hg is about 1/6 of that for Cs, due to the shorter wavelength of the allowed electric dipole moment transition of Hg, 180 nm compared to 895 nm and 854 nm (D1 and D2 transitions) for Cs, together with the reduced dielectric response of SiO2 at shorter wavelengths, leading to a reduction in the overlap integral.
+The value for Hg is about 1/6 of that for Cs, due to the shorter wavelength of the allowed electric dipole moment transition of Hg, 180 nm compared to 895 nm and 854 nm (D1 and D2 transitions) for Cs, together with the reduced dielectric response of SiO2 at shorter wavelengths.
+This leads to a reduction in the overlap integral.
 The number of bound states is approximately the number of nodes for the least tightly bound WKB solution, which is $n_{max}\approx 72$ for $m=199$, and the two lowest states are approximately -0.18 $e$V and -0.21 $e$V.
@@
 $\tau_{s}\propto\sum_{n=1}^{N}\tau_{0n}e^{|E_{an}|/kT}$ (8)
-where we allow $\tau_{0}$ to depend on $n$, and $E_{n}$ are the bound state energies.
+where we allow $\tau_{0}$ to depend on $n$, and $E_{n}$ are the bound state energies determined by the WKB approximation.
 However, since each level is likely to have a different sticking or residence time, a more profitable approach is to calculate the fraction of atoms on the surface by using the partition function.
-However, there is no need to consider a sticking time at this point.
-The binding energies of the levels are occupied according to the Boltzmann distribution, so we can determine the average fraction of the total number of atoms distributed among the bound states.
+The binding energies of the levels are occupied according to the Boltzmann distribution, so we can determine the average fraction of the total number of atoms distributed among the bound states and scale that according to the time between wall collisions.
 Assume that the atoms move unhindered while in the main cell volume (away from the wall) and at the wall there are bound states with energy $-E_{n}$ (for the following, we take $E_{n}=|E_{n}|$).
-While in a bound state, an atom can move freely parallel to the surface.
-Let the volume and area of, for simplicity, a spherical cell be $V$ and $A$, $\beta=1/k_{b}T$, and $p$ be the momentum of an atom.
-Using simple statistical theory, the probability $\epsilon_{n}$ that an atom spends time in the surface state $n$ is (assuming that this probability is much lower than 1 ),
+In a bound-surface state, we can assume that an atom moves more or less freely parallel to the surface.
+The model generally assumed for this problem is that the atom hops between lightly bound regions.[3] For simplicity, assume a spherical cell with volume and area $V$ and $A$, $\beta=1/k_{b}T$, and $p$ is the momentum of the atom.
+Using simple statistical theory, the probability $\epsilon_{n}$ that an atom spends time in the surface state $n$ is (assuming that this probability is dynamic and very small),
 $\epsilon_{n}$ $=$ $\frac{A\hbar^{-2}\int_{0}^{\sqrt{2mE_{n}}}\exp\left[\left(E_{n}-p^{2}/2m\right)\beta\right]d^{2}p}{V\hbar^{-3}\int_{0}^{\infty}\exp\left(-\beta p^{2}/2m\right)d^{3}p}$ (9)
@@
 where $\lambda_{T}=\hbar(2\pi\beta/m)^{1/2}$ is the thermal wavelength, $\beta=1/k_{b}T$, $m$ is the mass of an atom, and $R$ the radius of the cell, about 1 cm.
-For Hg at $293\mathrm{~{}K}$, $\lambda_{T}=$ $7\times 10^{-10}\mathrm{~{}cm}$ and using $E_{n}=0.21\ e$V with $R=1\mathrm{~{}cm}$ then $\epsilon=3\times 10^{-5}$.
+Approximately, for Hg at $293\mathrm{~{}K}$, $\lambda_{T}=$ $7\times 10^{-10}\mathrm{~{}cm}$.
+For the most strongly bound WKB state, $E_{n}\approx 0.21\ e$V, and with $R=1\mathrm{~{}cm}$ determines $\epsilon\approx 3\times 10^{-5}$.
 The total fraction of (temporarily) bound atoms on the surface is given by the sum of $\epsilon_{n}$ for all states.
@@
 The bound states as determined by the WKB approximation are shown in Fig.
-4, with an estimate of their cumulative probability distribution, the derivative of which gives the density of states.
+3, with an estimate of their cumulative probability distribution, the derivative of which gives the density of states.
 The total fraction of atoms on the wall is therefore
 […]
-and the integral is the “exponential integral function," ${\rm Ei}(E_{max}/k_{b}T)-{\rm Ei}(E_{min}/k_{b}T)$, and using Eq. (5.1.10) of 43,
+and the integral is the “exponential integral function," ${\rm Ei}(E_{max}/k_{b}T)-{\rm Ei}(E_{min}/k_{b}T)$, and using Eq. (5.1.10) of [54],
 $\epsilon\approx 3\pi\frac{\lambda_{T}}{R}\frac{n_{max}}{\log(E_{max}/E_{min})}\sum_{j=1}^{\infty}\frac{(E_{max}/k_{b}T)^{j}}{j\cdot j!}.$ (12)
 […]
-The relaxation rate should depend on the time that an atom is stuck on the surface; however, it is unclear if that is the product of a sticking time and a Boltzmann factor.
-Given that the time between wall collisions is $\approx v/R$ which is proportional to $\sqrt{T}$, then the relaxation rate should be
+The spin polarization relaxation rate should depend on the time that an atom resides on the surface; however, it is not clear whether $\tau_{w}$ is the product of a sticking time and a Boltzmann factor as has often been assumed.
+[3, 47] Given that the time between wall collisions is $\approx v/R$, which is proportional to $\sqrt{T}$, the relaxation rate should be
 $\Gamma_{w}\propto\sqrt{T}\epsilon\propto T\sum_{j=1}^{\infty}\frac{(E_{max}/k_{b}T)^{j}}{j\cdot j!}$ (13)
-because $\lambda_{T}\propto 1/\sqrt{T}$. Noting that the derivative of the summation in our original integrand,
-$\frac{1}{\Gamma_{p}}\frac{\partial\Gamma_{p}}{\partial T}=\frac{1}{T}-\frac{{e^{E_{max}/k_{b}T}}}{T}\left[{\sum_{j=1}^{\infty}\frac{(E_{max}/k_{b}T)^{j}}{j\cdot j!}}\right]^{-1}.$ (21)
-From the graph Fig.
-4 of [38], this ratio is about 0.03 at 260 K and 0.02 at 300 K, compared to the formula above numerically determined using $E_{max}=0.21$, with dielectric susceptibility the only specific material properties employed.
-Of course, correlations times, etc.
-will depend on temperature, but only weakly.
-Note that the collision rate times the sticking time $\Gamma_{p}\propto\sqrt{t}e^{\beta E_{0}}$ can be used to fit the data and approximately determines $E_{0}=0.15\ e$V which is close to $0.16\ e$V as obtained in [38] from the data presented in Fig.
-4 of that paper.
-Note that we did not need to consider a sticking time at all.
-There is another time scale in the problem, which is the free flight time between wall collision, which for a 1 cm dimension cell is about $10^{-4}$ s.
-If each traverse across the cell results in a sticking collision, then the time at atom is stuck in the potential wells should be $10^{-4}\epsilon=6\times 10^{-9}$ s to maintain the statistically necessary Boltzmann distribution of atoms.
-The correlation times in [38] are slightly shorter than this, which we might expect since the atoms move along the surface, the only requirement is that the sticking time be longer than the correlation time; the data in [38] implies the opposite, that the correlation time was less than the sticking time and was left as an open question.
-If the probability of a wall collision has a probability less than one for an atom to stick, the dwell time becomes longer, as required by microscopic reversibility and the value of $\epsilon$.
-This implies that the approximately 6 ns time obtained above is strictly a lower limit.
-The spatial extent of the surface binding region is only about 0.5 nm, so in $6\times 10^{-9}$ s a 199Hg atom will suffer $10^{4}$ collisions with the wall.
-Given that there are about 100 surface states, a random walk out of the miasma could very well require $\sim 100^{2}$ steps.
-The long wall sticking time for 129Xe as reported in [40] can be explained by there being only a small probability of capture in the wall potential for each wall collision, unlike the case of 199Hg discussed above, where it seems to be of order 50%.
-Given the chemical reactivity of Hg compared to Xe, this is not surprising, so a sticking probability for Xe between 0.1 and 1 % is needed to explain the results in [40].
-The reported binding energy of 0.1 $e$V as determined from the variation of the relaxation rate temperature is probably greater, as in the case of 199Hg where, without the full theory, 0.16 eV was obtained in [38] compared to $\approx 0.21$ eV for the most tightly bound surface obtained from the van der Waals interaction together with the WKB approximation.
+because $\lambda_{T}\propto 1/\sqrt{T}$. Noting that the fractional change in relaxation rate is readily found by taking the derivative of Eq. (12),
+$\frac{1}{\Gamma_{w}}\frac{\partial\Gamma_{w}}{\partial T}=\frac{1}{T}-\frac{{e^{E_{max}/k_{b}T}}}{T}\left[{\sum_{j=1}^{\infty}\frac{(E_{max}/k_{b}T)^{j}}{j\cdot j!}}\right]^{-1},$ (14)
+and therefore a direct comparison can be made with the data presented in Fig.
+4 of [47].
+The experimental values in [47] are 0.03 at 260 K and 0.02 at 300 K, compared to the numerically determined values 0.033 and 0.024 from Eq.
+(14), respectively, using $E_{max}=0.21$ from the WKB approximation.
+This agreement is quite good, determined by the use of only the frequency-dependent dielectric properties of fused silica together with other measured cell parameters (temperature, distances); in some sense the calculation is ab initio.
+Of course, correlations times will depend on temperature but only as the velocity.
+As an aside, it is of course possible to use the wall collision rate times the exponentially temperature dependent “sticking time," with $\Gamma_{w}\propto\sqrt{T}\tau_{0}e^{\beta E_{0}}$ in a fit to the data and approximately determine the “surface binding energy." The result is $E_{0}=0.15\ e$V which is close to $0.16\ e$V as obtained in [47] from the same data presented in Fig.
+4 of that article.
+Then, to get the sticking time, a value of $\tau_{0}$ is required.
+In [47], this is somewhat arbitrarily chosen as the inverse vibrational frequency of the CH4-Hg van der Waals complex, $\tau_{0}=10^{-12}$ s [55].
+The predicted sticking time at 293 K based on [47] is therefore $0.510^{-12}\ {\rm s}\ \times e^{0.16/.025}=$ 0.5 ns, which is approaching an order of magnitude lower than the observed correlation times of 2 ns.
+Alternatively, it is also possible to use the WKB results to estimate this frequency, by analogy with a harmonic oscillator and taking the difference in energy of the two most strongly bound vibrational states.
+This is about 0.02 eV, corresponding to a period $2\times 10^{-13}$ s.
+This could be expected because the molecular cluster vibrational frequency is determined by the reduced mass, which is essentially the CH4 mass; hence, the WKB spectrum has a lower energy by $\sqrt{16/199}=0.28$, which brings the frequencies into agreement.
+Using the WKB results in an even larger problem, a sticking time of 0.06 ns.
+It makes no sense for the correlation time to be less than the sticking time, so this approach does not really work.
+However, we do not need to consider a sticking time at all, but a dwell time.
+There is another time scale in the problem, which is the free flight time between wall collision, and for a 1 cm dimension cell, it is about $10^{-4}$ s.
+If each traverse across the cell results in a “sticking" collision, then the time an atom dwells in the potential well should be $10^{-4}\ {\rm s}\times\epsilon=6\times 10^{-9}$ s to maintain the statistically necessary Boltzmann distribution of atoms.
+The correlation times in [47] are slightly shorter than this, which might be expected because the atoms move along the surface while in a bound state and possibly react coherently only briefly with a given relaxing site.
+The only requirement is that the dwell time be longer than the correlation time.
+The data as interpreted in [47] implies the opposite, that the correlation time is longer than the sticking time, and this impossibility was left as an open question.
+If the probability that a wall collision results in an atom sticking is very small, the dwell time necessarily becomes longer, as required by microscopic reversibility and the value of $\epsilon$.
+This implies that the approximately 6 ns dwell time obtained above is strictly a lower limit.
+The spatial extent of the surface binding region is only about 0.5 nm, so in $6\times 10^{-9}$ s, while oscillating in the potential well, an 199Hg atom will suffer $10^{4}$ collisions with the wall.
+Given that there are about 100 surface states, a random walk out of the miasma of surface states could very well require $\sim 100^{2}$ steps, with $100^{2}\times 6\times 10^{-9}$ s $=6\times 10^{-5}$ s, which can be used to estimate the probability of wall collision sticking.
+For $10^{-4}$ s free flight time between wall collisions, the probability is $0.66\times 10^{-5}\ {\rm s}/10^{-4}\ {\rm s}=$ 0.6.
+The long wall sticking time for 129Xe as reported in [50] can be explained if there is only a small probability of capture in the wall potential for each wall collision, unlike the case of 199Hg discussed above, where it is found to be of order 60%.
+Given the chemical reactivity of Hg in comparison to Xe, this is not surprising, so the sticking probability for Xe is likely smaller, and the interaction with relaxing sites might be reduced.
+Both effects are needed to explain the results in [50].
+The reported binding energy of 0.1 $e$V, determined from the variation of the relaxation rate as a function of temperature, is probably too small.
+In the case of 199Hg where, without the full theory, 0.16 eV was obtained in [47] compared to $\approx 0.21$ eV for the most tightly bound surface obtained from the van der Waals interaction together with the WKB approximation, and the same conclusions about a “true" larger binding energy can be drawn in the case of Xe.
 Note that $\epsilon\propto 1/R$, whereas the free flight collision time is proportional to $R$, and therefore the overall relaxation rate is independent of $R$, as expected.
 We can conclude that instead of the characteristic vibration time on the wall, the relevant time scale is $\lambda_{T}/\bar{v}\propto\hbar/k_{b}T$.
-The idea of using the vibrational period of an atom near the wall as a sort of sticking time might be applicable for atoms with a bound state spectrum levels below $k_{b}T=0.025\ e$V.
-This idea was used in [44] from [45], and appears to not obviously apply to 199Hg or 129Xe.
+The possibility of using the vibrational period of an atom near the wall as a sort of sticking time is not without merit.
+For example, the vibrational period was used in the analysis of the relaxation of the electron spin polarization of alkali atoms as described in [56], based on the surface physics considerations discussed in [57].
 I.5.1 Hydrogen surface states
@@
 Also, the hard-sphere radius of H is smaller than that of Hg, so we can take $d_{0}\approx 0.12$ nm.
-Together with its smaller mass, this leads to the result that H has only four bound states at 0.08, 0.03, 0.01, and 0.003 $e$V.
-Although the energies are smaller than for Hg, the $\epsilon$ is enhanced by a relatively large $\lambda_{T}$ compared to Hg.
-The net result is $\epsilon=2\times 10^{-6}$, about an order of magnitude smaller than for Hg.
-Nevertheless, this result implies a microscopically long time for an excited Hg atom to interact with H atoms temporarily stuck to the surface.
-We can estimate the effective pressure of H on the surface, assuming that a fraction $62\times 10^{-}$ 6 of H atoms are imprisoned in a region about 0.3 nm thick.
-If there is 1 mbar H gas in a cell, the number density in the thin region near the cell wall corresponds to a pressure of about 4 bar.
-Alternatively, with a pressure $P_{0}$ in cell volume, the pressure near the surface (just like the isothermal atmosphere) should be $P_{0}e^{E_{max}/k_{b}T}=P_{s}$, and for $P_{0}=0.001$ bar, with an approximate binding energy of 0.21 $e$ V, it leads to $P_{s}=4.4$ bar and is astonishingly close to the previous estimate of 4 bar.
-Relaxation on the wall can be due to an enhanced interaction with H atoms or with the open bonds created when a surface is attacked by atomic H or excited Hg atoms.
+Together with the smaller mass of H, this leads to the result that H has only four bound states at 0.08, 0.03, 0.01, and 0.003 $e$V.
+Although the energies are smaller than for Hg, $\epsilon$ is enhanced by a relatively large $\lambda_{T}$ compared to Hg.
+The net result for $T=293K$, is $\epsilon=2\times 10^{-6}$, about an order of magnitude smaller than for Hg.
+Nevertheless, this result implies a microscopically long residence time for atomic H on the surface, during which Hg atoms can interact with the H electronic magnetic moment.
+We can estimate the effective pressure of H in the bound surface states, assuming that a fraction $\epsilon$ of the total H atoms are imprisoned in a region approximately $\delta=$ 0.3 nm thick.
+For a cell of radius $R$, the ratio of this surface volume to the cell volume is $3\delta/R$, so the relative densities are $\epsilon R/3\delta\approx 22$.
+For a given pressure $P$ of H gas in a 1 cm cell, the number density in the thin region near the cell wall corresponds to a pressure of about $22P$.
+Alternatively, with a pressure $P_{0}$ in cell volume, the pressure near the surface (similar to the isothermal atmosphere problem) should be $P_{0}e^{E_{max}/k_{b}T}=P_{s}$.
+For the maximum binding energy of 0.08 $e$ V for H from the WKB approximation, $e^{.08/.025}=24.5$ for $T=293$ K, which is in accord with the volume ratio determination.
+Spin-polarization relaxation on the wall can be due to an enhanced interaction with H atoms or with the open bonds created when a surface is attacked by atomic H or excited Hg atoms.
 These open (covalent) bonds carry roughly an electron magnetic moment.
@@
 I.6 Relaxation in the Gas phase
-Paramagnetic atoms and molecules, which include oxygen and free radicals, can exist in the buffer gas as impurities or the result of 254 nm radiation.
-Subsequent reactions with excited H can release atomic H (an unpaired spin, so technically a free radical) that exists in the states $F=0$ and $F=1$, the latter that exists as 3/4 of the total H and presents an effective 1/2 of an electron spin due to the addition of the proton nuclear angular momentum.
-The relaxation of nuclear spin $I=1/2$ 1S0 atoms by gaseous paramagnetic atoms is theoretically analyzed by Jameson, Jameson, and Hwang [46] for the specific case of O2 and 129Xe, and further developed in [47], for both O2 on 3He and on 129Xe Rewritten here in terms of SI units, and more generally for a colliding molecule (or atom) $X$ with total angular momentum $F$ representing the sum of the internal angular momentum of the constituent nuclei and electron spins of $X$,
+Paramagnetic atoms and molecules, which include oxygen and free radicals, can exist in the buffer gas as impurities or the result of a 254 nm radiation induced photochemical reaction.
+Reactions with an excited Hg atom can release atomic H (an unpaired spin, which is technically a free radical) that exists in the states $F=0$ and $F=1$, in proportion to 1/4 and 3/4 of total H and $F+1$ presents an effective 1/2 of an electron magnetic moment due to the addition of the proton nuclear angular momentum.
+The relaxation of nuclear spin-polarized $I=1/2$ 1 S0 atoms by gaseous paramagnetic atoms is theoretically analyzed by Jameson, Jameson, and Hwang [58] for the specific case of O2 and 129Xe, and further developed in [59], for O2 on both 3He and 129Xe.
+Rewritten here in terms of SI units and more generally for a colliding molecule (or atom) $X$ with total angular momentum $F$ representing the sum of the internal angular momentum of the constituent nuclei and the electron spins of $X$,
 $\frac{1}{T_{1}}=\frac{16}{3}F(F+1)~{}\gamma_{I}^{2}~{}\left[\mu_{0}~{}g_{F}~{}\gamma_{e}\right]^{2}\frac{\hbar^{2}}{d^{2}}\left(\frac{\pi\mu}{8k_{B}T}\right)^{1/2}~{}F(U,T)~{}n_{0}\left[X\right]$ (15)
-$\gamma_{I}$ and $\gamma_{e}$ are the gyromagnetic ratios for the $I=1/2$ nucleus, and for the bare electron respectively; $g_{F}$ is the Landè $g-$ factor for $X$, and $\mu_{0}=4\times 10^{-7}$.
+$\gamma_{I}$ and $\gamma_{e}$ are the gyromagnetic ratios for the $I=1/2$ nucleus, and for the bare electron respectively; $g_{F}$ is the Landè $g-$ factor for $X$, and $\mu_{0}=4\pi\times 10^{-7}$.
 $d$ is the characteristic intermolecular interaction length, $\mu$ is the reduced mass of the colliding species, $k_{B}$ is the Boltzmann constant and $T$ is the absolute temperature (the factor in parentheses is the reciprocal of the mean relative velocity).
@@
 The factor $F(U,T)$ is determined by the interaction potential $U$ between the colliding species, and is unity at high temperature in the hard-core limit.
-In general, it is determined by integrating solutions of the radial Schr"odinger equation with the potential $U$, together with an average over the Boltzmann momentum distribution.
-It is noted in [47] that the hyperfine Fermi contact term, proportional to $|\psi(0)|^{4}$, should generally be included in $F(V,T)$, where $\psi(0)$ is the amplitude of the electron wave function of the colliding molecule at the nucleus $I$.
+In general, it is determined by integrating solutions of the radial Schrödinger equation with the potential $U$, together with an average over the Boltzmann momentum distribution.
+In [59] it is noted that the hyperfine Fermi contact term, proportional to $|\psi(0)|^{4}$, should generally be included in $F(U,T)$, where $\psi(0)$ is the amplitude of the electron wave function of the colliding molecule at the nucleus $I$.
 The interaction time (correlation time) for gas phase molecules is very short, so effects due to an applied magnetic field become small when the electron precession over the collision time results in a small directional change in the electron(s) spin angle(s); therefore, in all cases under consideration such effects can be ignored for the gas phase.
-In [47], the experimental relaxation rate for 3He in O2 is reported to be approximately $0.5$ [s Amagat], which implies $F(U,T)=1.7$.
-For 129Xe, $F(U,T)\approx 6$ is estimated to be [46] which implies a relaxation rate of 0.7/[s Amagat].
-Data in [48] indicates that the relaxation rate of 199Hg in O2 is 800/[s Amagat], which about 1000 times faster than for the inert gases 3He and 129Xe.
-Using the analysis for inert gases implies that $F(U,T)=1.5\times 10^{4}$ although the two valence electrons of 199Hg likely enhance a hyperfine interaction between the atom and the molecule, the analysis for inert gases is not adequate.
-Another possibility is that van der Waals molecules can be formed, in which case the needed three-body collisions would make the relaxation rate pressure dependent.
+In [59], the experimental relaxation rate for 3He in O2 is reported to be approximately $0.5$ /(s Amagat), which implies $F(U,T)=1.7$.
+For 129Xe, $F(U,T)\approx 6$ is estimated in [58], which implies a relaxation rate of 0.7/(s Amagat).
+(An Amagat is the number density of an ideal gas at a pressure of 1 bar or 760 torr at 273.16 K.)
+Data in [60] indicate that the spin polarization relaxation rate of 199Hg in O2 is 800/(s Amagat), which is approximately 1000 times faster than for the inert gases 3 He and 129 Xe.
+Using the analysis for inert gases implies that $F(U,T)=1.5\times 10^{4}$ for 199Hg .
+Because the two valence electrons of 199Hg enhance a hyperfine interaction between the atom and the molecule, the analysis for inert gases is not entirely accurate for Hg.
+Another possibility is that van der Waals molecules between 199Hg and O2 can be formed, in which case the needed three-body collisions would make the relaxation rate pressure dependence more complicated.
 This would be interesting to study.
-We can try to apply this result to the relaxation of 199Hg by H.
-Because H has a nuclear spin of 1/2, 3/4 of the atoms will be in the hyperfine level with $F=1$ and $\gamma_{s}$ is reduced by a factor of two compared to O2.
-Comparing $F(U,T)$ with previous cases of inert gases suggests that $F(U,T)$ scales as the square root of the reduced mass, which we might expect from the Schrödinger equation.
-Scaling by $\mu$ and $\gamma_{s}^{2}$, and by the fraction of atoms in $F=1$ leads to a relaxation rate of approximately 5/[s Amagat].
-The general conclusion is that the pressure of O2 in a room-temperature cell must be below 6 mbar to achieve a spin lifetime of 200 s.
-In the case of Fig.
-1, to achieve a lifetime of 3600 s requires a pressure of 0.3 mbar.
-This implies that the purity of the gas used to fill the cell to 233 mbar must have been about 99.9%, which is the typical level of commercial gases of non-UHP grade.
-A direct measure of the H relaxation rate would require the production of atomic H and a means to measure its pressure.
-Based on the foregoing, a total H2 pressure less than a few mbar appears as desirable in a cell, particularly given the destructive and autocatalytic behavior of atomic H.
+The general conclusion is that the pressure of O2 in a room temperature cell must be less than the 6 $\mu$bar level to achieve $\tau_{w}>200$ s.
+This implies that the oxygen impurity of the nitrogen gas used to fill the cell in Fig.
+1 was below 10 parts per million, consistent with its spectroscopic purity grade.
+Even with this high purity, the oxygen concentration in cells made using the same nitrogen had a rapid loss of Hg vapor under illumination with resonance radiation (the Hg filling pressure was 1.5 $\mu$ bars corresponding to the 0 vapor pressure, much lower than the oxygen pressure), as illustrated in Fig.
+2.
+We can apply the foregoing analysis to the spin polarization relaxation of 199Hg by atomic H.
+Because H has a nuclear spin of 1/2, 3/4 of the atoms will be at hyperfine level with $F=1$ and $\gamma_{s}$ is reduced by a factor of two compared to O2.
+Comparing $F(U,T)$ with previous cases of inert gases suggests that $F(U,T)$ scales as the square root of the reduced mass of the 199Hg -H system, as expected from the Schrödinger equation.
+Scaling by $\mu$ and $\gamma_{s}^{2}$, and by the fraction of H atoms in $F=1$ leads to a relaxation rate for 199Hg nuclear spin polarization of approximately 5/(s Amagat).
+A direct measure of the H-induced spin polarization relaxation rate would require the production of atomic H and a means to measure its pressure.
+In light of such measurements, based on the foregoing, a total H2 pressure well below a few millibars appears to be necessary to stabilize $\tau_{w}$ in a cell, particularly given the destructive and autocatalytic behavior of atomic H.
 II Materials and Methods
@@
 4 was assembled in Fall 1991 at the University of Washington (these data are being reported for the first time).
-The system consisted of a selection of gases (99 + % purity) that could be introduced into an optical pumping cell contained in a three-layer magnetic shield, together with vacuum pumps and optical pumping components.
+The system consisted of an optical pumping cell in a three-layer magnetic shield and optical components.
+The cell was connected to a vacuum system through a manifold that provided a selection of high purity gases (99.99% or better) or air that could be introduced into the cell.
 The cell itself was of volume 750 $\ell$ (we refer to these as “1 liter cells") based on a design used to evaluate the optical pumping of 199Hg in larger volumes than had ever been done before, and was a key R&D component toward the development of the system described in [20].
+Figure 4: Manifold and optical system used for relaxation tests. See text for details.
 The cell was coated with a generic pure PTFE lubricating spray in a hydrocarbon carrier.
-Cells with a coating made using a pure PTFE lubricating spray in a primarily hydro/fluorocarbon carrier (with minor components of isopropyl alcohol and acetone, without water) had better stability than other fluorinated coatings or any water-based Teflon coatings.
-An example is Elmer’s Slide-All® (which is no longer manufactured); although it was manufactured in a fast-evaporating water-free carrier, it contained a more complex polymer that has hydrogenous functional groups on the ends of PTFE polymer chains, and these groups are subject to attack by excited-state Hg or atomic H.
+Cells with a coating made using a pure PTFE lubricating spray in a primarily hydro/fluorocarbon carrier (with minor components of isopropyl alcohol and acetone, without water) had better stability than other partially fluorinated coatings or any water-based Teflon coatings.
+An example is Elmer’s Slide-All® (which is no longer manufactured); although it was manufactured with a fast-evaporating water-free carrier, it contained a more complex polymer with hydrogen-containing functional groups at the ends of the PTFE polymer chain, as shown in Fig.
+5.
+These functional groups are subject to attack by excited-state Hg or atomic H and showed less stability than pure PTFE.
 A lubricating spray that uses the same molecule is still available, MicroCare® VDX Dry Lubricant Spray, Aerosol.
@@
 Using Eq.
-1, the mean free path is
+2, the mean free path is
 $\ell=\frac{4V}{A}=\frac{4\pi R^{2}h}{2\pi R^{2}+2\pi Rh}=\frac{2h}{1+h/R}=6.7\ {\rm cm}$ (16)
-where $R$ is the inner radius, and $h$ is the height.
-These coatings typically gave about 20 s of lifetime per cm mfp, which implies a lifetime of about 130 s for this large cell.
+where $R$ is the inner radius and $h$ is the height.
+These coatings typically gave about 20 s of $\tau_{w}$ per cm mfp, which implies a $\tau_{w}$ of about 130 s for this large cell.
 The 1 mm fill hole, if complete depolarization occurs for 199Hg atoms entering that region, will lead to a relaxation rate of
 $\Gamma_{hole}=\frac{1}{4}\frac{\bar{v}A_{h}}{V}\approx 40\ {\rm s}$ (17)
-where $\bar{v}$ is the average Hg velocity and $A_{h}$ is the hole area, as based on the kinetic theory of atomic effusion.
-The observed cell lifetime was about 136 s, implying that the 199Hg atoms had a good chance of returning to the main volume with their polarization.
+based on the atomic flux determined by kinetic theory.
+The observed cell $\tau_{w}$ was about 136 s, implying that the 199Hg atoms had a good chance of returning to the main volume with their polarization.
 This is because the capillary inside is coated with PTFE and which, given the smoothness of the surface, indicates that Hg atoms stick briefly to the surface and re-emitted in a random direction.
-Figure 6: Manifold and optical system used for relaxation tests. See text for details.
 The operation of the system was straightforward.
 Natural Hg was supplied to the cell from a droplet until the transmitted light was reduced by a factor of 2.
-Various gases could be admitted and their pressures measured using a Hg manometer or McLeod gauge.
+Various gases could be admitted, and their pressures could be measured using a Hg manometer or McLeod gauge.
 A magnetic field of about 10 mG was applied along the light direction to polarize the atoms, and could be suddenly switched perpendicular to the light to observe the modulated light transmission due to the precessing atoms, as measured with a solar-blind photomultiplier.
 The magnetic shields had a substantial internal magnetic gradient; these shields were constructed from scrap materials and used for testing.
-The gradient limited the $T_{2}$ lifetime with gases at higher pressures; however, in those cases the longitudinal polarization pump-up and decay could be observed, giving $T_{1}\approx\tau_{0}$, the wall relaxation time.
+The gradient limited $T_{2}$ $\tau_{w}$ with gases at higher pressures; however, in those cases, the longitudinal polarization pump-up and decay could be observed, giving $T_{1}\approx\tau_{0}$, the wall relaxation time.
 The pumping was carried out with the full light intensity, which gave a pumping time of 60 s, while the decay was observed with the full light and also with the light intensity reduced by a factor of 10 (using an optical attenuating filter), which allowed extracting the wall decay time $\tau_{w}$.
-The main goal of the work was to assess the effect of hydrogen on the wall coating used in [14] which degraded with the application of high voltage, the application of which, we surmised, created atomic H due to inevitable microdischarges on dielectric surfaces.
+The main goal of the work was to assess the effect of hydrogen on the wall coating used in [20] for which $\tau_{w}$ degraded with the application of high voltage (HV).
+We surmised that the application of HV created atomic H as a result of inevitable microdischarges on the dielectric surfaces or dirt in the system.
 Previous work (see Fig.
-1) suggested that oxygen might restore the cell relaxation lifetime.
-The secondary goal was to test for this possibility and determine how much oxygen could be used before limiting the 199Hg lifetime.
+1) suggested that oxygen might restore the cell $\tau_{w}$.
+The secondary goal was to test for this possibility and determine how much oxygen could be used before limiting the 199Hg spin polarization relaxation time.
 The tests were performed with PTFE instead of deuterated polystyrene (DPS) [21] in part because in 1991 the final principal edm cell coating material used in [20] had not yet been decided.
-Typically, dps provides lifetimes of about 15 s per cm mfp, about as good as PTFE.
+Typically, DPS provides $\tau_{w}$ of about 15 s per cm mfp, about as good as PTFE.
 Resources were not available to continue the 1991 studies with DPS.
@@
 Test 6 shows that hydrogen gas alone does not damage the wall coating.
-Tests 7-12 show that Hg in the presence of resonance light immediately degrades the cell, and removal of the gas does not repair the degraded lifetime.
-Given its stability, the degradation is probably not due to the atomic H being adsorbed on the surface but is more likely due to unterminated bonds on the wall coating.
+Tests 7-12 show that Hg in the presence of resonance light immediately degrades the cell, and removal of the gas does not repair the degraded $\tau_{w}$.
+Given its stability, the degradation is probably not due entirely to the fact that atomic H is adsorbed on the surface but also to unterminated bonds on the wall coating.
 Atomic H has enough energy to do this damage, as shown in Table 1, and reported in [46, 45].
@@
 Tests 15 and 17 indicate that the damage might be local as the buffer gas can tend to hide damage.
-The 254 nm light entered at exited the cell in a beam of about 2 cm diamter, so it is conceivable that the most significant coating damage was in these areas.
-Table 2: Tests conducted with the apparatus shown in Fig. 6. All times have an uncertainty of 8 s unless noted otherwise.
+The 254 nm light entered and exited the cell in a beam of about 2 cm diameter, so it is conceivable that the most significant coating damage was in these areas.
+Table 2: Tests conducted with the apparatus shown in Fig. 4. All $\tau_{w}$ measurements have an uncertainty of 8 s unless noted otherwise.
 Test No. Condition $\tau_{w}$ [s] Comment/Result
-1 only Hg vapor 136 Cell starting lifetime
+1 only Hg vapor 136 Cell starting $\tau_{w}$
 2 add 1.5 torr N2, 5% CO 133
@@
 14 increase to 50 torr " no signal
-15 increase to 110 torr " 145±$\pm 30$ pump up method so low signal; improvement
-with buffer gas implies localized damage
+15 increase to 110 torr " 145±$\pm 30$ pump up method so low signal;
+improvement with buffer gas
+implies localized damage
 16 evacuate, add Hg 36 partial recovery?
-17 add 110 torr N2, 5%CO $\approx 60-100$ similar to 15
+17 add 110 torr N2, 5%CO $80\pm 20$ similar to 15
 18 evacuate, add Hg 36 no apparent further recovery
@@
 20 evacuate, add Hg 121 oxygen exposure nearly recovered $\tau_{w}$
-21 add 0.65 torr air $\approx 7$ relaxation rate $\approx 1100$/[s Amagat]
-22 evacuate, add Hg 134 sat. Hg density, then evacuate to 1/2 abs.
-23 evacuated, add Hg 110 evacuation was overnight $2\times 10^{-7}$ torr
-24 1 hr irradiation 167 lifetime improving on irradiation
+21 add 0.65 torr air $\approx 7$ O2 relaxation rate $\approx$840/(s Amagat)
+22 evacuate, add Hg 134 fill to high Hg density,
+then evacuate to 1/2 light absorption
+23 evacuate, add Hg 110 evacuation overnight, to $2\times 10^{-7}$ torr
+24 1 hr irradiation 167 $\tau_{w}$ improving on irradiation
 25 no change - stabilized at this value
 IV Discussion and Conclusions
-Based on the literature on photosensitized Hg reactions, together with our own observations, the results clearly indicate that hydrogen, through atomic H, is nearly certainly responsible for the degradation of 199Hg optically pumped magnetometry cells.
+Based on the literature on photosensitized Hg reactions, together with our own observations, the results clearly indicate that hydrogen, through atomic H, is nearly certainly responsible for much of the degradation of $\tau_{w}$ in 199Hg optically pumped magnetometry cells.
 Atomic H arises from the reaction of H2 with excited state Hg.
@@
 Our 1991 measurement of the relaxation of 199Hg by oxygen gas agrees well with the most recent measurement reported in [60].
-Given the salubrious effect of oxygen on cell lifetimes, the inclusion of a small amount will be beneficial; however, it will be consumed to form substances that stick to the wall or mix with wax.
-In addition, Hg will be consumed, so a source of 199Hg is needed within the cell, or ozone production must be suppressed.
-Given that of order $10^{13}$ photon per second are incident on a cell and that the photosensitize reaction in some cases has a quantum efficiency near unity, we might expect the same number of atoms or molecules lost per second.
-This corresponds to $10^{18}$ reaction per day, or 10 mbar of gas produced per day in a 5 $\ell$ cell.
+Given the salubrious effect of oxygen $\tau_{w}$, the inclusion of a small amount will be beneficial; however, it will be consumed to form substances that stick to the wall or mix with wall coating materials.
+In addition, Hg will be consumed to form HgO, so a source of 199Hg within the cell is needed, or ozone production must be suppressed.
+Given that of order $10^{13}$ photon per second are incident on a cell and that the photosensitized reaction in some cases has a quantum efficiency near unity, we might expect the same number of atoms or molecules lost per second.
+This corresponds to $10^{18}$ reaction per day, or 10 mbar of gas consumed per day in a 5 $\ell$ cell.
 According to Table 1, the only (single) bond that is immune to damage is HF.
@@
 Hydrogen likely does its damage through an autocatalytic reaction, in that if there is any hydrogen, it will lead to photosensitized production of more hydrogen by breaking the bonds of hydrogenous materials on the cell walls.
-However, fluorinated waxes appear to not work as well as hydrocarbon waxes, but more studies are needed.
 It was hoped that CO might bind to atomic H and eventually form formaldehyde, which would get buried in the wall coating.
@@
 Identifying the surface states allows for a theoretically consisted picture to explain the data presented in [47]. In all, this was quite remarkable and unexpected.
-Similarly, the seemingly anonymously long sticking times for 129Xe as reported in [40] can be explained by such states together with a low probability of sticking per wall collision, which is expected for Xe compared to Hg.
-Using microscopic reversibility and the calculable value of $\epsilon$ as the fraction of atoms stuck in the wall potential at any given time, if the dwell time can be calculated, perhaps using the Boltzmann transport equation for motion among the myriad of surface states, then the sticking probability is determined.
-Oddly enough, the reduced interaction of Xe probably means it stays on the surface longer, as the heating interactions that will lead to its release are likely smaller than those for Hg.
-Regardless, there is another time scale that is better suited for the case of heavy atoms with many bound states, and that is the thermal wavelength divided by the average velocity, which of course is proportional $\hbar$ divided atom kinetic energy, which instead of the thermal wavelength, is the inverse of the thermal frequency.
-An important consequence of the large number of relatively deep surface states (compared to $k_{b}T=0.025$ eV) is in regard to the assumption that a precise volume magnetic field average is provided by 199Hg co-magnetometry.
-For an extremely precise magnetometry system, as has been proposed, the surface dwell time (during which a substantial spin polarization survives) will need to be more carefully considered, as the dwell time can lead to corrections.
-For example, localized systematically generated magnetic fields are sensitive to the volume averaging.
+Similarly, the seemingly anonymously long sticking times for 129Xe as reported in [50] can be explained through surface states together with a low probability of sticking per wall collision, which is expected for Xe compared to Hg.
+Using microscopic reversibility and the calculable value of $\epsilon$ as the average fraction of atoms stuck in the wall potential, the probability of sticking can be determined, which is of the order 1%.
+Rather than the sticking time, there is another time scale that is better suited for the case of heavy atoms with many bound states.
+That time scale is the thermal wavelength divided by the average velocity, which is proportional $\hbar$ divided the kinetic energy of the atom.
+This time scale is the inverse of the “thermal frequency,” in analogy to the thermal wavelength.
+The large number of relatively deep surface states (compared to $k_{b}T=0.025$ eV) brings into question the efficacy of 199Hg co-magnetometry.
+For an extremely precise magnetometry system, as has been proposed, the surface dwell time (during which a substantial spin polarization can survive) will need to be more carefully considered, as the dwell time can lead to corrections, e.g., localized systematically generated magnetic fields are sensitive to the volume averaging.
 The gravitation offset of ultracold neutrons in neutron EDM experiments is perhaps a more serious problem; however, all such effects will eventually need consideration.
-Based on the foregoing, to improve the durability of closed or open 199Hg optical pumping and detection cells (i.e., any similar cells used for fundamental experiments), some guidelines are as follows:
+Based on the foregoing, to improve the durability of closed or open 199Hg optical pumping and detection cells (i.e. any similar cells used for fundamental experiments), some guidelines are as follows:
   1. 1.
-    Avoid hydrogen gas either as an impurity in the CO buffer gas (oxygen can be removed, but water as an impurity will produce hydrogen which will pass through the amalgam) or generated by photosensitized reactions.
+    Avoid hydrogen gas either as an impurity in the CO buffer gas (oxygen can be removed, but water as an impurity will produce hydrogen which will pass through the purification amalgam) or generated by photosensitized reactions.
 Sealing of filled fused silica cells involves much heat and even UV light is produced, so care is needed in such operations;
   2. 2.
-    For open cells used in the neutron EDM experiment that employs a 199Hg comagnetometer, the problem of lifetime degradation appears to be largely solved.[16] This is not surprising as the benzene ring is relatively immune from attack by atomic H [49], and exposure to ozone and oxygen lead to crosslinkages in the polymer.[50]
+    For open cells used in a neutron EDM experiment that employs a 199Hg comagnetometer, the problem of $\tau_{w}$ degradation appears to be largely solved, in particular by use of an oxygen gas rinse.
+[23, 22] The use of deuterated polystyrene offers some stability from attack by atomic H and other free radicals.
+This is in part due to the 1.6 eV energy required to disrupt the resonance stabilization of the phenyl ring.
+With this additional energy, an H atom cannot cause a C-H bond to break to form H-H; however, atomic H can be added to the phenyl ring, thus producing a cyclohexadienyl radical,[61] while exposure to ozone and oxygen leads to cross-linkages in the polymer.[62] This additional energy presents a barrier to direct reactions between benzene molecules and excited-state Hg atoms, as can be seen from the energies in Table 1.
+A study of the reaction between atomic O and benzene used the photosensitized reaction between Hg and N2O to produce O in situ in a reaction vessel, with the final amount of nitrogen as a measure of the total amount of reacted oxygen.[63]
+    Pertaining to high voltage stability, the phenyl ring is attacked by positive ions that can bind to the electron cloud of the ring, leading to destabilization of the molecule due to disturbance of the resonance of the ring.
+This implies that DPS-coated cells require that high voltage can be applied without electrical discharges that produce positive ions.
+This also implies that DPS is less stable under high voltages than aliphatic compounds.
   3. 3.
-    Photosensitized reactions can be suppressed by use of a quenching gas; CO has proven effective but there are better gases;
+    Photosensitized reactions in closed cells can be suppressed by use of a quenching gas; CO has proven effective but there are better gases;
   4. 4.
@@
   6. 6.
-    Fluorinated waxes or silanes with less hydrogen, e.g., Trichloro(3,3,3-trifluoropropyl)silane (CAS 592-09-6).
-A fully fluorinated silane would be of interest; however, none appear to be commercially available.
+    Use fluorinated waxes or silanes with less hydrogen, e.g., Trichloro(3,3,3-trifluoropropyl)silane (CAS 592-09-6).
+A fully fluorinated silane would be of interest; however, none appears to be commercially available.
+Tests with fluorinated wax have not shown significant improvement but appear as promising.[43, 65]
   7. 7.
-    Deuterated polystyrene[15] remains one of the few options for UCN 199Hg comagnetometer experiments.
+    Deuterated polystyrene (DPS) [21] remains one of the few options for UCN 199Hg comagnetometer experiments.
 Acknowledgments
@@
 All data used in this study are presented in this report or are fully contained in the cited literature.
+S.K.L. is solely responsible for the conception, research, writing, and finalization of this article. The author will read and agree to the published version of the manuscript.